### v1

Abstract
Time series anomaly detection (TSAD) finds many applications such as monitoring environmental sensors, industry KPIs, patient biomarkers, etc. A two-fold challenge for TSAD is a versatile and unsupervised model that can detect various different types of time series anomalies (spikes, discontinuities, trend shifts, etc.) without any labeled data. Modern neural networks have outstanding ability in modeling complex time series. Self-supervised models in particular tackle unsupervised TSAD by transforming the input via various augmentations to create pseudo anomalies for training. However, their performance is sensitive to the choice of augmentation, which is hard to choose in practice, while there exists no effort in the literature on data augmentation tuning for TSAD without labels. Our work aims to fill this gap. We introduce TSAP for TSA “on autoPilot”, which can (self-)tune augmentation hyperparameters end-to-end. It stands on two key components: a differentiable augmentation architecture and an unsupervised validation loss to effectively assess the alignment between augmentation type and anomaly type. Case studies show TSAP’s ability to effectively select the (discrete) augmentation type and associated (continuous) hyperparameters. In turn, it outperforms established baselines, including SOTA self-supervised models, on diverse TSAD tasks exhibiting different anomaly types.
1 Introduction
Anomaly detection (AD) is a critical task in various domains such as cybersecurity, healthcare, and finance. AD is especially important in time series to ensure system safety and reliability. Thus, there exists a large body of work on time series AD as presented in various surveys [5, 15]. Recent progress on self-supervised learning (SSL) has changed the field of AD, offering significant improvements over traditional unsupervised (or one-class) learning approaches. The main advantage of SSL lies in its ability to self-generate labeled samples, i.e., pseudo anomalies, within the input space. This enables a more focused exploration of a plausible subspace based on the semantics reflected in the pseudo anomalies, rather than an exhaustive, impractical search of the entire space.
At the heart of SSL-based AD are data augmentation functions. These functions are used to create pseudo labels for the (self-)supervised training of an anomaly detector, such as by predicting whether the input is augmented or not [20], which augmentation function is used [14], or using contrastive learning [31]. In all these approaches, the success of SSL-based AD highly depends on the degree to which the augmented data mimics the true anomalies, as outlined in [35].
There exist few approaches for tuning data augmentation functions without labels, however, they have limitations. Some rely on non-differentiable validation losses [36], which are unsuitable for end-to-end learning frameworks. Others only address continuous hyperparameters, neglecting discrete ones, which are left for manual adjustment [34]. For instance, in the image domain, the CutOut augmentation cannot mimic semantic class anomalies (e.g. cats vs. cars), no matter how one tunes its (continuous) hyperparameters [34], mainly due to the mismatch in the discrete choice (i.e., the augmentation type). Moreover, none of these efforts addresses anomaly detection for time series data, which is the focus of our work.
In this paper we introduce TSAP, a novel approach for SSL-based time series anomaly detection (TSAD) “on autoPilot” equipped with end-to-end hyperparameter tuning. Effectively tuning both discrete and continuous hyperparameters of augmentation enables TSAP to be an automated anomaly detector that is most suitable for a given task. TSAP stands on two main components: ($i$) a differentiable parameterized augmentation model and ($ii$) an unsupervised validation loss that measures the alignment between the augmented data and the unlabeled test data, quantifying the extent to which the former mimics the true anomalies. We summarize our main contributions as follows:
  1. 1.
    Problem: Our work is the first attempt to tune both discrete and continuous hyperparameters of data augmentation in SSL-based TSAD, without any labels at training time.
  2. 2.
    New TSAD Method: We propose TSAP  [^1] The code will be published once the paper is accepted., which accommodates a list of various time series anomaly types and enables automatic tuning of related hyperparameters (magnitude, duration, etc.) by a differentiable validation loss that quantifies alignment of augmented data and unlabeled test data.
  3. 3.
    Effectiveness: By carefully selecting augmentation type and its (continuous) hyperparameters, our self-tuning TSAP outperforms existing unsupervised and self-supervised approaches, including the SOTA NeuTraL-AD [26] which also employs learnable augmentations, across diverse TSAD tasks.
2 Preliminaries
2.1 Time Series Anomaly Detection
This work focuses on time series anomaly detection. Consider a univariate time series $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{K}\}$, where $\mathbf{x}$ is a sequentially ordered collection of $K$ data points. Each $x_{k}\in\mathbb{R}$ corresponds to a scalar observation at each time step. Then, let $\mathcal{D}_{\mathrm{trn}}$ be a set of training data consisting of only normal time series, and $\mathcal{D}_{\mathrm{test}}$ be a set of unlabeled test data containing both normal and anomalous time series. Given $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{test}}$, the problem is to detect all anomalous time series in $\mathcal{D}_{\mathrm{test}}$, i.e., to assign the accurate label $y\in\{-1,+1\}$ for each $\mathbf{x}\in\mathcal{D}_{\mathrm{test}}$.
2.2 Self-supervised Anomaly Detectors
Let $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ be a data augmentation function, such as rotation or cut-out in the image domain (a discrete choice), that outputs pseudo anomalies. A given $f_{\mathrm{aug}}$ also exhibits continuous hyperparameter(s), such as angle (for rotation) or width and height (for cutout). Then, a popular approach [20] is to train a detector model $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta})$ to classify between normal data $\mathcal{D}_{\mathrm{trn}}$ and augmented data $\mathcal{D}_{\mathrm{aug}}=\{f_{\mathrm{aug}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$. The working assumption is for $f_{\mathrm{det}}$ to predict the (unknown) anomalies in test data as augmented. As such, performance relies on how well the choice of augmentation type (discrete) as well as the choice(s) of its (continuous) hyperparameter(s) mimic the anomalies in $\mathcal{D}_{\mathrm{test}}$. There are various other ways of using SSL for anomaly detection, based on the specific objective function and how the pseudo labels are generated [14, 31], which are similar to each other in principle.
2.3 Data Augmentation on Time Series
To apply SSL to time series data, we need an augmentation function specifically designed for time series. For a given time series $\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}$, $f_{\mathrm{aug}}(\mathbf{x};\mathbf{a})$ transforms $\mathbf{x}$ based on hyperparameters $\mathbf{a}\in\mathcal{A}^{P}$, $P\geq 1$, where $\mathcal{A}^{P}$ represents the domain of all possible hyperparameter values. For time series data, augmentations could involve mean shifts, trend injections, or other transformations relevant to the time-series domain, with $\mathcal{A}^{P}$ encompassing the respective hyperparameter space for a given transformation. We describe augmentation in more detail in Sec. 3.1.
2.4 Wasserstein Distance
The Wasserstein distance [1], a distance measure between probability distributions, of order $p$ between any two marginal distributions $\mu$ and $\rho$ is given as
$W_{p}(\mu,\nu)=\biggl{(}\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}[d(x,y)^{p}]\biggr{)}^{1/p},$ (1)
[…]
3 TSAP: Time Series Anomalies on AutoPilot
Figure 1: Our TSAP framework for end-to-end self-tuning TSAD. Left: Offline trained, differentiable augmentation model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ takes as input the normal data and augmentation hyperparameter(s) $\mathbf{a}$, and outputs pseudo-anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$. Right: Self-tuning engine incorporates the pre-trained $f_{\mathrm{aug}}$ (with parameters $\boldsymbol{\phi}$ frozen), alternating between two phases: ($i$) detection phase – given $\mathbf{a}^{(t)}$ at iteration $t$, estimate parameters $\boldsymbol{\theta}^{(t)}$ of detector $f_{\mathrm{det}}$ (consisting of Encoderθ and discriminator MLPθ), by optimizing $\mathcal{L}_{\mathrm{trn}}$ (classification loss); ($ii$) alignment phase – given $f_{\mathrm{det}}^{\mathrm{enc}}(\cdot;\boldsymbol{\theta}^{(t)})$, update augmentation (governed by $\mathbf{a}$) to better align the embedded time series $\mathbf{z}_{\mathrm{trn}}\cup\mathbf{z}_{\mathrm{aug}}$ with $\mathbf{z}_{\mathrm{val}}$ in the learned discriminative space. Note that $\mathbf{x}_{\mathrm{val}}$ contains both normal and anomalous time series, but labels are not known or used at any point during training time.
There are two notable challenges that need to be addressed for automatic selection of both discrete and continuous hyperparameters for SSL-based TSAD:
Challenges
  • C1.
    Differentiable Augmentation: Developing an augmentation function that is differentiable with respect to its hyperparameters, enabling gradient-based optimization.
  • C2.
    Comparable Validation Loss: Formulating a validation loss that effectively quantifies alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{test}}$ while being comparable across different hyperparameter initializations.
To address C1 and C2, we center our framework around two main ideas:
Main ideas
  • I1.
    Differentiable Augmentation Module: TSAP implements augmentation function as an Encoder-Decoder neural network, $f_{\mathrm{aug}}$ parameterized by $\boldsymbol{\phi}$, capable of approximating the anomaly-generating mechanism conditioned on $\mathbf{a}\in\mathcal{A}^{P}$. Importantly, this module is pre-trained independently and prior to the initiation of I2, establishing it as an offline component of the framework.
  • I2.
    Self-Tuning Module: At test time online, TSAP iteratively refines the detector $f_{\mathrm{det}}$’s parameters $\boldsymbol{\theta}$ as well as augmentation hyperparameters $\mathbf{a}$, through alternating detection and alignment phases. Alignment is performed on part of the unlabeled $\mathcal{D}_{\mathrm{test}}$, referred to as $\mathcal{D}_{\mathrm{val}}$.
Based on the key ideas I1 and I2, we propose TSAP, a self-tuning self-supervised TSAD framework illustrated in Fig. 1.
Figure 2: Examples of six different types of time series anomalies; (black) original real-world time series, (red) pseudo anomalies generated by $g$. The goal for $f_{\mathrm{aug}}$ is to learn to inject anomalies into inliers, conditional on the discrete (anomaly type) as well as continuous hyperparameters such as level (magnitude), location (start point) and length (see bottom right). More details are presented in Sec. 3.1 and Appx. 0.A.1.
3.1 Differentiable Augmentation Module
Anomaly Injection Scheme
We carefully consider accommodating six types of anomalies that are common in real-world time series; namely, trend, extremum, amplitude, mean shift, frequency shift, and platform. We provide a detailed description of the anomalies in Appx. 0.A.1 and for brevity, illustrate them by examples in Fig. 2 with red lines. Each type of anomaly has three hyperparameters; including its starting position (location), duration (length), and severity (level) as illustrated in Fig. 2 (bottom right). Extremum captures a spike and only has two hyperparameters as its duration is always $1$. Based on $\mathcal{D}_{\mathrm{trn}}$, the anomaly generation scheme $g$ creates an augmented dataset $\mathcal{D}_{\mathrm{aug}}=\{g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})\mid\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}},\mathbf{a}\sim\mathcal{A}^{P}\}$, where $\mathbf{a}$ is the vector of augmentation hyperparameters that are uniformly randomly sampled from the domain $\mathcal{A}^{P}$.
Model Design
To build an augmentation model $f_{\mathrm{aug}}$ for time series, first we use a convolutional neural network (CNN) to encode or embed the input time series $\mathbf{x}_{\mathrm{trn}}$ into the feature map $\mathbf{z}_{\mathrm{trn}}$. We then encode the augmentation hyperparameters $\mathbf{a}$ into $\mathbf{z}_{\mathrm{a}}$, which has the same shape as $\mathbf{z}_{\mathrm{trn}}$, with a multilayer perceptron (MLP). Since the feature map $\mathbf{z}_{\mathrm{trn}}$ generated by the CNN encoder keeps the positional information of the original time series, adding $\mathbf{z}_{\mathrm{a}}$ to $\mathbf{z}_{\mathrm{trn}}$ ensures that only the part with the desired location and length in $\mathbf{z}_{\mathrm{aug}}$ is manipulated. To ensure that the feature maps $\mathbf{z}_{\mathrm{aug}}$ and $\mathbf{z}_{\mathrm{trn}}$ are in the same embedding space, they share the same decoder to reconstruct back to the time series $\mathbf{x}_{\mathrm{aug}}=g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ and $\mathbf{x}_{\mathrm{trn}}$, respectively. As such, the loss function of $f_{\mathrm{aug}}$ is based on the reconstruction of both $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{aug}}$:
$\mathcal{L}_{\mathrm{aug}}=\sum_{\begin{subarray}{c}\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}}\\ \mathbf{a}\sim\mathcal{A}^{P}\end{subarray}}\big{(}{(\mathbf{x}_{\mathrm{trn}}-\mathbf{\tilde{x}}_{\mathrm{trn}})}^{2}+{(g(\mathbf{x}_{\mathrm{trn}},\mathbf{a})-\mathbf{\tilde{x}}_{\mathrm{aug}})}^{2}\big{)}$ (2)
where $\mathbf{\tilde{x}}_{\mathrm{aug}}$ denotes the output of $f_{\mathrm{aug}}(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ for the same value of the hyperparameters sampled by $g$ (see Fig. 1 left).
Algorithm 1 Self-Tuning Module of TSAP
1:Input: Train & val. data $\mathcal{D}_{\mathrm{trn}}$& $\mathcal{D}_{\mathrm{val}}$, pre-trained aug. model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$, detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta})$, max epoch $T$, #inner-loops $L$, and step sizes $\alpha$ and $\beta$
2:Output: Optimized aug. hyperparameters $\mathbf{a}^{*}$ and optimized parameters $\boldsymbol{\theta}^{*}$ corresponding to the lowest $\mathcal{L}_{\mathrm{val}}^{*}$
3:$\mathbf{a}^{(0)}\sim\mathrm{Uniform}(\mathcal{A}^{P})$
4:for $t=0,1,\ldots,T-1$ do
5:    $\boldsymbol{\theta}^{\prime}\leftarrow\boldsymbol{\theta}-\alpha\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\mathrm{trn}}(\boldsymbol{\theta},\mathbf{a}^{(t)})$ for $l$ in $L$triangleright Phase ($i$)
6:    $\mathcal{Z}_{\mathrm{trn}}\leftarrow\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$triangleright Phase ($ii$) starts
7:    $\mathcal{Z}_{\mathrm{aug}}\leftarrow\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x},\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$
8:    $\mathcal{Z}_{\mathrm{val}}\leftarrow\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{val}}\}$
9:    $\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\leftarrow\mathrm{norm.}(\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}})$
10:    $\mathbf{a}^{(t+1)}\leftarrow\mathbf{a}^{(t)}-\beta\nabla_{\mathbf{a}}\mathcal{L}_{\mathrm{val}}(\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}})$
11:    $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}^{\prime}$triangleright Phase ($ii$) ends
12:    if $\mathcal{L}_{\mathrm{val}}^{(t)}<\mathcal{L}_{\mathrm{val}}^{*}$ then $\mathbf{a}^{*}\leftarrow\mathbf{a}^{(t+1)}$;   $\boldsymbol{\theta}^{*}\leftarrow\boldsymbol{\theta}$ end if
13:end for
14:Return $\mathbf{a}^{*}$ and $\boldsymbol{\theta}^{*}$
3.2 Self-Tuning Module
Central to TSAP is the self-tuning module, which operates by iteratively refining the detector’s parameters, $\boldsymbol{\theta}$, and the augmentation hyperparameters, $\mathbf{a}$. The process is structured into two phases: detection and alignment (see Fig. 1 right). The overall algorithm is given in Algo. 1.
Phase ($i$): Detection
This phase focuses on estimating the parameters $\boldsymbol{\theta}^{(t)}$ of the detector $f_{\mathrm{det}}$ (comprising of an encoder $f_{\mathrm{det}}^{\mathrm{enc}}$ and a discriminator $f_{\mathrm{det}}^{\mathrm{mlp}}$) by minimizing the cross-entropy loss $\mathcal{L}_{\mathrm{trn}}$. This aims to classify between the normal samples $\mathbf{x}_{\mathrm{trn}}$ and the augmented pseudo anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$ by their embeddings $\mathcal{Z}_{\mathrm{trn}}$ and $\mathcal{Z}_{\mathrm{aug}}$, where $\mathcal{Z}_{\mathrm{trn}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ and $\mathcal{Z}_{\mathrm{aug}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x};\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ denote the embeddings of the training data and augmented data, respectively, given the current fixed $\mathbf{a}^{(t)}$ at iteration $t$. Note that the parameters $\boldsymbol{\phi}$ of the augmentation model $f_{\mathrm{aug}}$ are frozen throughout this phase.
Phase ($ii$): Alignment
Subsequently, the alignment phase adjusts $\mathbf{a}$ to optimize the unsupervised differentiable validation loss $\mathcal{L}_{\mathrm{val}}$, computed based on the embeddings from the now-updated $f_{\mathrm{det}}^{\mathrm{enc}}$. $\mathcal{L}_{\mathrm{val}}$’s objective is to measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space, as expressed by the Wasserstein distance in Eq. (1). This metric is chosen for its effectiveness in capturing the overall distributional discrepancies between datasets, offering a more complete comparison than mere point-wise metrics [1], which is especially important in TSAD given the often subtle nature of time series anomalies. Note that the embeddings are normalized to ensure scale invariance before being passed to $\mathcal{L}_{\mathrm{val}}$, so as to avoid the trivial solution of achieving alignment by setting all embeddings to the zero vector [34]. Lastly, as the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ in $\mathcal{L}_{\mathrm{val}}$ are obtained through the updated $f_{\mathrm{det}}^{\mathrm{enc}}$, the optimization process needs to track the change in $\boldsymbol{\theta}$ caused by the update of $\mathbf{a}$. As such, TSAP employs a second-order optimization process, similar to [34].
Augmentation Type Selection
While Algo. 1 describes continuous hyperparameter tuning, the discrete hyperparameter (i.e. augmentation type) selection, corresponding to a specific anomaly type, is done through grid search as the number of anomaly types is finite. Hence, we initialize TSAP for different augmentation types and compare $\mathcal{L}_{\mathrm{val}}$ across types to select the one that yields the best alignment. The idea is that the wrong augmentation type will have poor alignment, while one that reflects the true anomalies in $\mathcal{D}_{\mathrm{test}}$ will result in better alignment, granted proper tuning of the continuous hyperparameters through Algo. 1.
4 Experiments
In this section, we aim to answer the following questions:
  1. Q1.
    Quantitative Comparison: How well does TSAP perform on TSAD against established baselines?
  2. Q2.
    Qualitative Analysis: How well does TSAP tune the augmentation hyperparameters?
  3. Q3.
    Ablation Studies: Which design decisions in TSAP drive its effectiveness?
4.1 Setup
Datasets
We evaluate TSAP on six distinct TSAD tasks. Four of these are conducted in a controlled environment, while the remaining two are natural. In the controlled setting, the anomaly types are manually injected in $\mathcal{D}_{\mathrm{test}}$ based on the types discussed in Sec. 3.1. The controlled environment allows for a thorough assessment of TSAP’s ability to perform continuous hyperparameter tuning. For the natural environment, the anomaly types in $\mathcal{D}_{\mathrm{test}}$ are a priori unknown and it is the goal of TSAP to find the anomaly type that yields best alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ (part of $\mathcal{D}_{\mathrm{test}}$), expressed by $\mathcal{L}_{\mathrm{val}}$. The details of the datasets and anomaly types are provided in Appx. 0.A.1 and 0.A.2.
For controlled tasks, we use the 2017 PhysioNet Challenge dataset [9], which comprises real-world ECG recordings. Table 1 shows a selection of four different controlled TSAD tasks, constructed by manually injecting different anomaly profiles into the PhysioNet dataset. For PhysioNet A and B, given the anomaly type (Platform), the task is to infer or tune, respectively, the hyperparameter(s) level only and both level and length, while anomaly location is random (hence not tuned). For PhysioNet C and D, the respective tuning tasks are the same but for a different anomaly type (Trend). We choose Platform and Trend anomalies to represent, respectively, range-bound and range-exceeding anomalies. Additional variations of PhysioNet with other injected anomaly types, such as Mean shift and Extremum, are discussed in Appx. 0.A.2.
The natural TSAD tasks are derived from the CMU Motion Capture (MoCap) dataset [^2] http://mocap.cs.cmu.edu/. We consider the walking signal as normal data, and the signals of jumping and running as anomalies. To generate normal signals, we stitch the walking signals by identifying the start and end points of each gait phase and add random noise; whereas to generate anomalous ones, we stitch walking or running signals at a random location in the normal signal. This yields two distinct TSAD tasks as shown in Table 1. Different from PhysioNet A–D where we only tune the continuous hyperparameter(s) for the given (discrete) anomaly type, here for MoCap A and B, we aim to tune both the unknown anomaly type that corresponds to Jump and Run behavior, respectively, as well as the (continuous) hyperparameter level while location and length take random values.
Table 1: Anomaly profile of different TSAD tasks.
Profile PhysioNet A PhysioNet B PhysioNet C PhysioNet D MoCap A MoCap B
Type Platform Platform Trend Trend Jump Run
Level Fixed Fixed Fixed Fixed Fixed Fixed
Location Random Random Random Random Random Random
Length Random Fixed Random Fixed Random Random
Baselines
We compare TSAP with a selection of established baselines, including both traditional and deep learning-based methods that have demonstrated efficacy in TSAD [29]. Among the traditional methods, we utilize three with different modeling approaches; namely, One-Class Support Vector Machines (OC-SVM) [30]; Local Outlier Factor (LOF) [7]; and Isolation Forest (IF) [21]. On the deep learning side, we benchmark against the Encoder-Decoder LSTM (EncDec-LSTM) [24]; the Spectral Residual Convolutional Neural Network (SR-CNN) [28]; and the Unsupervised Anomaly Detection (USAD) for TSAD [2]. Lastly, we include a state-of-the-art competing method which learns augmentations in the embedding space, called Neural Transformation Learning for (TS)AD (NeuTraL-AD) [26]. This diverse set of baselines allows for a comprehensive analysis across different approaches within the TSAD domain.
Model Configurations
The Encoderϕ in $f_{\mathrm{aug}}$ and Encoderθ in $f_{\mathrm{det}}$ are constructed using 1D CNN blocks [4] (transposed 1D CNN for Decoderϕ) for efficient temporal feature extraction. We carefully choose the number of epochs $T$ to allow sufficient time for the convergence of $\mathbf{a}$, with empirical evidence suggesting that $T=100$ typically suffices. For the number of inner-loops $L$, we opt for $L=5$, aligned with [34], such that $f_{\mathrm{det}}$ has adequate time to learn effective discriminative embeddings for $\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{trn}}$. A detailed overview of the configuration for TSAP and all baselines is provided in Appx. 0.A.3.
Evaluation Metrics
Our method calculates anomaly scores on an entire sequence level $\mathbf{x}$, similar to [26]. This is a different set-up compared to novelty detection in time series which typically operates on a point level. Detection on a sequence level can be especially important to spot long-term patterns (e.g. Trend anomalies). As such, we use the Area Under the Receiver Operating Characteristic Curve (AUROC) and the $F_{1}$ score as key performance metrics to quantify detection capability of anomalous sequences. All results are reported on the unseen $\mathcal{D}_{\mathrm{test}}$. In addition, we determine the optimal $F_{1}$ score, by enumerating all possible thresholds, given by the anomaly scores for a given segment $\mathbf{x}$. We then compute the corresponding precision and recall for each threshold and select those that yield the highest $F_{1}$. As such, AUROC provides a balanced view whereas $F_{1}$ shows the optimal scenario. Both metrics range from 0 to 1, with higher values indicating superior performance.
4.2 Quantitative Results
Table 2: Detection peformance of baselines w.r.t. $F_{1}$ and AUROC on test data across six TSAD tasks. Four tasks are performed in a controlled setting (anomaly type is manually injected and corresponds to anomaly types described in Sec. 3.1), based on the PhysioNet ECG data. Remaining two TSAD tasks exhibit natural anomalies (real-world anomaly type which is unknown), based on the MoCap data. TSAP outperforms most baselines and has the best (lowest) average rank w.r.t. both $F_{1}$ and AUROC, with low standard deviation (in parentheses).
Methods PhysioNet A PhysioNet B PhysioNet C PhysioNet D MoCap A MoCap B Avg. Rank
$F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC
OC-SVM 0.182 0.468 0.182 0.472 0.373 0.803 0.393 0.806 1.000 1.000 0.546 0.806 5.7 (2.6) 5.3 (2.3)
LOF 0.999 1.000 0.999 1.000 0.354 0.738 0.358 0.725 0.196 0.506 0.221 0.577 5.5 (2.7) 5.2 (3.3)
IF 0.255 0.587 0.232 0.576 0.183 0.402 0.182 0.356 0.864 0.965 0.342 0.758 6.2 (1.6) 6.3 (1.4)
EncDec-LSTM 0.190 0.508 0.190 0.508 0.415 0.812 0.442 0.819 0.980 0.999 0.909 0.996 4.8 (1.0) 5.00 (0.9)
SR-CNN 0.965 0.990 0.964 0.998 0.983 0.999 0.991 0.999 0.302 0.700 0.214 0.512 3.8 (3.0) 4.0 (2.8)
USAD 0.183 0.425 0.184 0.428 0.430 0.822 0.409 0.828 1.000 1.000 1.000 1.000 4.2 (2.7) 4.3 (3.1)
NeuTraL-AD 0.211 0.732 0.263 0.679 0.561 0.868 0.526 0.862 1.000 1.000 1.000 1.000 2.8 (1.6) 2.7 (1.4)
TSAP (ours) 1.000 1.000 1.000 1.000 0.973 0.999 0.991 0.998 0.889 0.969 1.000 1.000 1.8 (1.6) 1.8 (1.6)
Table 2 provides the detection results for all six TSAD tasks described in Sec. 4.1. TSAP ranks the best overall in terms of both $F_{1}$ and AUROC. This shows that detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$, trained through the alternating mechanism of TSAP, is able to generalize to unseen and unlabeled anomalies in $\mathcal{D}_{\mathrm{test}}$.
While some competing methods perform strongly on a subset of tasks, they lack consistency across all TSAD tasks. Among the traditional baselines, which perform subpar as compared to deep TSAD approaches, LOF performs very well in PhysioNet A and B, relatively poorly on C and D, but exhibits near-random performance on MoCap. This discrepancy can be attributed to the nature of the anomalies in the PhysioNet data, which are manually injected and often characterized by pronounced and abrupt changes as shown in Fig. 2. These abrupt changes lead to large differences in local density, which makes LOF thrive. In contrast, the MoCap dataset contains anomalies that emerge more subtly, as one gradually transitions from walking to running or jumping, reflecting more natural variations in the data which, in turn, lead to more subtle differences in local density. Similarly, SR-CNN performs strongly on the PhysioNet TSAD tasks, yet its effectiveness also diminishes drastically on MoCap tasks. SR-CNN exploits the abrupt changes present in the PhysioNet data but fails to detect the more subtle anomalies in MoCap. This suggests that, while SR-CNN and LOF work well for certain types of anomalies, their utility may be limited in scenarios where anomalies are more subtle.
On the contrary, reconstruction-based baselines such as EncDec-LSTM and USAD face challenges on PhysioNet due to its high variability among inliers, which complicates the task of accurate data reconstruction. Yet, these methods excel with the MoCap dataset, where its consistent near-periodic pattern (Fig. 2, bottom) lends itself to more reliable reconstruction, thereby enhancing anomaly detection performance. Similarly, NeuTraL-AD, a state-of-the-art augmentation-based baseline, demonstrates proficiency in properly augmenting the inlier data within MoCap. However, its performance on the PhysioNet variations reveals some weaknesses. NeuTraL-AD struggles with the high variability in the PhysioNet inliers, inherent to real-world ECG signals. This suggests that the augmentation functions considered in NeuTraL-AD lack robustness when inliers are inherently noisy.
We remark that only TSAP provides robust and consistent performance across all TSAD tasks, showcasing the effectiveness and generalizability of our proposed method.
4.3 Qualitative Results
Figure 3: (best in color) Tuning continuous augmentation hyperparameter(s) with TSAP. Top: Given Platform anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (center), and leading to high detection performance (right). Bottom: Multiple continuous hyperparameters, here both level and length, are accurately tuned to near true values (left), as guided by minimizing the val. loss (center), achieving high AUROC (right). Notice that the diverged results in both cases associate with high (or hard-to-optimize) val. loss, which help us effectively reject low performance models.
A key contributor to TSAP’s consistent performance is the validation loss through which TSAP automatically learns the augmentation hyperparameters $\mathbf{a}$. Once $\mathbf{a}$ is determined, the task essentially reduces to a supervised learning problem. Next, we show that TSAP not only effectively tunes the continuous augmentation hyperparameters $\mathbf{a}$, but also that its validation loss guides the accurate selection of the discrete hyperparameter (i.e. anomaly type).
Controlled Environment
Consider PhysioNet A, where we aim to tune the continuous hyperparameter $a$, i.e. level, of the Platform anomalies present in $\mathcal{D}_{\mathrm{test}}$. That is, the level in $\mathcal{D}_{\mathrm{test}}$ is fixed and tuning aims to estimate its value using TSAP while the other hyperparameters (location, length) are randomized. Fig. 3 (top) shows TSAP’s estimation process for different initializations of $a$. We observe that the initialization for $a\in\{-0.4,0.6\}$ leads to the true $a$$=$$0.2$ (left). Simultaneously, the validation loss drops substantially once TSAP has arrived at the true $a$ (middle). This is also reflected in the performance of $f_{\mathrm{det}}$ on part of the unlabeled test data $\mathcal{D}_{\mathrm{val}}$ which soars upon estimation of the true $a$ (right). Conversely, the initialization for $a\in\{0.8,-0.8\}$ leads to a high validation loss, indicating poor alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$. Indeed, the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ now suffers from poor alignment.
For PhysioNet B, we estimate both level and length while location is randomized. Fig. 3 (bottom) demonstrates TSAP’s ability to accurately estimate the level and length. While this is a by-product of our method, there are several real-world use-cases that can directly benefit from accurately learning the anomaly profile, from industrial equipment monitoring, to network security, and healthcare monitoring. This demonstrates TSAP’s versatility given its capability of estimating continuous hyperparameters. See Appx. 0.A.4 for additional qualitative results on other anomaly types such as Trend, Mean shift, and Spike.
Further, Fig. 4 (top) showcases TSAP’s ability to perform discrete hyperparameter selection. Here, TSAP has been initialized and trained with three different anomaly types (Mean shift, Platform, Trend) on PhysioNet C (true anomaly type is Trend). The validation loss clearly showcases a misalignment between the Platform and Trend types (left), which is also reflected in the AUROC of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ (middle). Note how the Mean shift anomaly type has a low $\mathcal{L}_{\mathrm{val}}$ towards the end of the training epochs, which is also reflected in the AUROC on $\mathcal{D}_{\mathrm{val}}$. Indeed, comparing TSAP tuned for Mean shift anomalies and for Trend anomalies show strong resemblances with the true underlying anomaly type (right). This demonstrates that the true underlying anomaly type is not necessarily the only type that yields high alignment, and in turn, a high-performing detector.
Natural Environment
In MoCap datasets, the anomaly types are a priori unknown. As such, we initialize TSAP with different augmentation types (Frequency and Platform) to perform discrete hyperparameter selection. Fig. 4 (bottom) highlights its effectiveness as the validation loss clearly prefers one type over the other. Indeed, the natural anomalies defined by jumping signals in MoCap A have close resemblance to platform anomalies. Discrete hyperparameter optimization for MoCap B is similar, as given in Appx. 0.A.4.
Figure 4: (best in color) Tuning the discrete hyperparameter (anomaly type) with TSAP. Top: Given true Trend anomaly (black), val. loss favors both Trend (green) and Mean shift (blue) type, both with high AUROC (middle) and high resemblance (right), and effectively rejects type Platform (orange) with poor performance. Bottom: For Jump anomalies in MoCap A with unknown type (black), val. loss favors type Platform (purple) that leads to high AUROC (middle) and mimics well the true anomaly (right), and effectively rejects type Frequency (red) with poor performance.
4.4 Ablation Studies
We showcase two ablation studies on the controlled PhysioNet data to support various design strategies of TSAP. Additional ablation studies on embedding normalization and second-order optimization are provided in Appx. 0.A.4.
Validation Loss
In Fig. 5 (top), we illustrate the level estimation for PhysioNet C under the condition where our Wasserstein-based $\mathcal{L}_{\mathrm{val}}$ is substituted with a point-wise metric, as used in [34]. This comparison shows that a point-wise validation loss tends to favor solutions where the level of the Trend anomaly approximates zero, essentially neutralizing the anomaly. Although this might produce high alignment, it leads to poor $f_{\mathrm{det}}$ performance in $\mathcal{D}_{\mathrm{val}}$ (right). This shows that the distributional characteristics captured by our $\mathcal{L}_{\mathrm{val}}$ are a key contributing factor to the success of TSAP.
Figure 5: (best in color) Overview of ablation studies. Top: TSAP’s $\mathcal{L}_{\mathrm{val}}$ is replaced with a point-wise val. loss leading to an erroneous estimation of $a$ (left) and poor performance (right). Bottom: TSAP’s self-tuning module is disabled and $a$ is instead randomized. Val. loss indicates poor alignment, reflected in poor performance.
Randomization vs. Tuning
In Fig. 5 (bottom), the self-tuning module is disabled for PhysioNet C, where level is instead randomized (along with location, and length). We observe substantially higher $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment. In turn, $f_{\mathrm{det}}$ struggles to detect the unlabeled anomalies in $\mathcal{D}_{\mathrm{val}}$. This clearly demonstrates the utility of TSAP’s systematic hyperparameter (self-) tuning over random choice.
5 Related Work
Self-supervised Learning (SSL) for Anomaly Detection:
SSL has emerged as a significant approach in machine learning. Foundation models [6] like large language models [25, 27], whose training heavily relies on SSL, have shaken up the world with outstanding performance. SSL has also transformed representation learning and significantly boosted several tasks in NLP, computer vision, recommender systems and medicine [3, 19]. SSL is especially attractive for unsupervised problems like anomaly detection (AD) [16], because of its nature to create proxy tasks and loss functions without any labels. There are various types of SSL-based methods on AD [22], but the core idea is to use data augmentation functions [8, 14, 20] on inliers to create pseudo anomalies, and learn a classifier that can detect the pseudo anomalies.
Time Series Anomaly Detection:
There exists a large body of work on TSAD, for which we refer to several surveys [5, 15]. Recently, SSL has been widely studied as the main approach to address TSAD problems [32, 37]. BeatGAN [43] and RobustTAD [13] use data augmentation to enrich training data. COUTA [33] proposes three different augmentations, respectively mimicking local, contextual and collective anomalies. NeuTraL-AD [26] propose several augmentation functions that are an integral part of the learning process. AMSL [38] also uses signal transformations as data augmentation. TimeAutoAD [18] employs three different strategies to augment the training data for generating pseudo anomalies. Motivated by masked language models [11], MAD [12] investigated various masking procedures for multivariateTSAD based on a predictive pretext task. Note that only [18, 26] parameterize the augmentation procedures, but they tune those using labeled validation data.
Unsupervised Model Selection:
Unsupervised hyperparameter tuning (i.e. model selection) is nontrivial in anomaly detection due to the absence of labeled validation data [23], and the literature is slim with quite recent efforts [39, 40, 41, 42]. Specifically on SSL-based AD, a recent study [35] has revealed the impact of the choice of augmentation on SSL-based AD, leading to several works that aim to automatically search for the optimal choice [34, 36]. Their main contribution is a novel unsupervised loss, but it is not differentiable. None of the existing works addressed the model selection problem of SSL-based AD for time series data.
6 Conclusion
We introduced TSAP for self-supervised time series anomaly detection, which is the first attempt that automatically (self-)tunes the augmentation hyperparameters on time series data in an unsupervised manner. TSAP includes a differentiable model to augment the input time series data with various anomaly types, and an unsupervised validation loss that assists in aligning the augmented and test data. Experiments showed TSAP’s ability in effectively selecting the augmentation type along with its continuous hyperparameters. Across various real-world datasets with different types of time series anomalies, TSAP outperformed a diverse list of baselines, including modern neural and self-supervised approaches. While being the first self-tuning SSL solution to TSAD, our work opens new research directions, including an expanded catalog of supported anomaly types and possible extension of the ideas to multivariate time series data.
Appendix 0.A Appendix
0.A.1 Types of Time Series Anomalies
We include six types of time series anomalies:
  • Platform: Starting at timestamp location, the values of a duration length in the time series are equal to a constant value level.
  • Mean shift: Starting at timestamp location, a constant value level is added to the values of a duration length in the time series.
  • Amplitude: Starting at timestamp location, a constant value level is multiplied with the values of a duration length in the time series.
  • Trend: Starting at timestamp location, a series of values $at$ is added to the duration length, where $a$ is the level and $t$ is the timestamp in that duration.
  • Extremum (a.k.a. Spike): A large (either positive or negative) value level is assigned to a single timestamp location in the time series.
  • Frequency shift: Starting at phase location, the frequency of the duration with length phases is increased by a constant value level.
Fig. 2 visualizes each type of anomaly and demonstrates how each hyperparameter controls the injection of the anomaly in the time series for platform anomaly.
0.A.2 Dataset Details
PhysioNet
The 2017 PhysioNet Challenge dataset [9] comprises a diverse array of real-world 1-lead 300 Hz ECG recordings. We use ECG recordings, each 9 seconds in length with $K=2700$ time-steps and standardized to have a zero mean and standard deviation of one. Injected anomalies represent 10% of the data. Here, we include three additional controlled TSAD tasks based on the PhysioNet data as shown in Table 3.
Table 4 (top) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in PhysioNet. The hyperparameters location and length are normalized by $K$. Noting that the extremum anomaly always occurs on a single timestamp in the time series, thus length is always $1$.
Table 3: Anomaly profile of additional TSAD tasks E–G.
Profile PhysioNet E PhysioNet F PhysioNet G
Type Mean shift Mean shift Extremum
Level Fixed Fixed Fixed
Location Random Random Random
Length Random Fixed N/A
MoCap
The CMU Motion Capture (MoCap) dataset [^3] http://mocap.cs.cmu.edu/ includes signal data from various sensors on subjects’ bodies as they perform different activities (walking, jumping, or running). As we focus on a univariate setting, only the sensor signal on the left femur is used. Since each data contains the time series data with a short length, we stitch them into a longer one in our experiment. To ensure it is done smoothly, we identified the start and end points of each gait phase for the stitching. We further add random noises to augment the normal samples in the dataset. Each signal is normalized between $-1$ and $1$ and truncated to length $K=1500$. Constructed anomalies represent 10% of the data. As opposed to PhysioNet, we consider MoCap as a dataset with natural or real anomalies. Therefore, we do not create additional TSAD tasks from this dataset.
Table 4 (bottom) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in MoCap. The hyperparameters location and length of platform anomaly are normalized by $K$. The hyperparameters location and length of frequency anomaly denote the starting gait phase and the length of gait phases, respectively.
Table 4: Hyperparameter space for each type of anomaly.
PhysioNet
Anomaly Types location length level
Platform $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Mean Shift $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Amplitude $[100..2000]$ $[400..600]$ $\{0.5k+1|k\in[0..10]\}$
Trend $[100..2000]$ $[400..600]$ $\{0.002k-0.01|k\in[0..10]\}$
Extremum/Spike $\{100k|k\in[1..26]\}$$\{1\}$ $\{3k-15|k\in[0..10]\}$
MoCap
Anomaly Types location length level
Platform $[200..800]$ $[100..200]$ $\{0.2k-1|k\in[0..10]\}$
Frequency Shift $[1..3]$ $[1..6]$ $[1..3]$
0.A.3 Model Configurations
Following Sec. 4.1, we provide further details on the configurations for TSAP and all baselines.
TSAP configuration:
In Table 5, we provide a comprehensive overview of the configuration details for the different components of TSAP. Table 5 (top) shows the configuration details for the different components of $f_{\mathrm{aug}}$ following Fig. 1 (left). Similarly, Table 5 (bottom) shows the configuration details for the components of $f_{\mathrm{det}}$ following Fig. 1 (right).
Table 5: Configuration details for $f_{\mathrm{aug}}$ and $f_{\mathrm{det}}$ of TSAP
$f_{\mathrm{aug}}$ configuration
Encoderϕ
   Conv Layer 1 {In: 1, Out: 64, Kernel: 100, Stride: 4, ReLU, BatchNorm}
   Conv Layer 2 {In: 64, Out: 64, Kernel: 100, Stride: 4, ReLU}
Decoderϕ
   TransConv Layer 1 {In: 64, Out: 64, Kernel: 100, Stride: 4, ReLU, BatchNorm}
   TransConv Layer 2 {In: 64, Out: 1, Kernel: 100, Stride: 4}
MLPϕ
   MLP Layer 1 {In: 3, Out: 16, ReLU}
   MLP Layer 2 {In: 16, Out: dim$(\mathcal{Z}_{\mathrm{trn}})$, ReLU}
General Parameters
   Batch Size 64
   # Epochs 500
   Optimizer Adam (LR: 0.002)
$f_{\mathrm{det}}$ configuration
Encoderθ
   Conv Layer 1 {In: 1, Out: 32, Kernel: 10, Stride: 2, ReLU, BatchNorm}
   Conv Layer 2 {In: 32, Out: 16, Kernel: 10, Dilation: 2, Stride: 2}
   Conv Layer 3 {In: 16, Out: 8, Kernel: 10, Dilation: 4, Stride: 4}
   Avg Pooling + Flatten {Kernel: 10, Stride: 3}
   Linear Layer {In: 400, Out: 10}
MLPθ
   MLP Layer 1 {In: 10, Out: 1}
General Parameters
   Dropout 0.2
   Batch Size 64
   Warm Start # Epochs $f_{\mathrm{det}}$ 3
   # Epochs 100
   Optimizer for $\mathbf{a}$ Adam (LR: 0.001)
   Optimizer for $f_{\mathrm{det}}$ Adam (LR: 0.002)
   Mixing Rate (cf. Phase ii) 0.15
Baseline configurations:
We used the parameter settings recommended by the original authors for all baseline models. This approach ensures that our comparisons are fair and align with the intended set-up of these methods originally proposed by the authors.
  • OC-SVM: We use author-recommended hyperparameters based on [17].
  • LOF: We use author-recommended hyperparameters based on [7].
  • IF: We use author-recommended hyperparameters based on [21].
  • EncDec-LSTM: Similar to the original authors [24], we downsample our time series to obtain time series of length approx. equal to 200. For the remaining hyperparameters, we use the original authors’ recommendations.
  • SR-CNN: We use author-recommended hyperparameters based on [28].
  • USAD: Similar to the original authors [2], we downsample our time series to obtain time series of length approx. equal to 200. For the remaining hyperparameters, we use the original authors’ recommendations.
  • NeuTraL-AD: We use author-recommended hyperparameters based on [26]. We tune augmentation type separately for each dataset using labeled validation data.
0.A.4 Additional Results
We present additional results for continuous and discrete augmentation hyperparameter tuning, as well as two additional ablation studies.
0.A.4.1 A.4.1 Continuous Augmentation Hyperparameter Tuning
In addition to the results on continuous hyperparameter tuning for PhysioNet A and B (see Fig. 3), we demonstrate TSAP’s efficacy in tuning the continuous hyperparameters on five additional TSAD tasks. These include PhysioNet C and D, which feature Trend anomalies (cf. Table 1), as well as PhysioNet E and F, showcasing Mean shift anomalies, and PhysioNet G, illustrating Extremum anomalies (cf. Table 3).
PhysioNet C & D:
Figure 6: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Trend anomalies. Top: Given Trend anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (middle), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, here both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (middle), achieving high AUROC (right).
The tuning process of the continuous hyperparameters for the Trend anomalies in PhysioNet C and D is shown in Fig. 6. We observe several initializations for $\mathbf{a}$ that arrive closely to the true level for PhysioNet C, as well as to the true level and length for PhysioNet D. In turn, those initializations yield a detector $f_{\mathrm{det}}$ with high performance on $\mathcal{D}_{\mathrm{val}}$. Note how, for PhysioNet C, $\mathcal{L}_{\mathrm{val}}$ is low across the board. This is likely due to the fact that a Trend anomaly with a subtle slope, has similar characteristics to inliers (see e.g. Fig. 4 bottom left). Yet, TSAP effectively assigns a higher validation loss to the initializations that lead to misaligned cases. This shows the effectiveness of our method even in cases where anomalies are subtle.
PhysioNet E & F:
Similarly to Trend anomalies, Mean shift anomalies are inherently subtle, especially when the level is close to zero. To this end, we show in Fig. 7 how TSAP properly tunes the continuous hyperparameters for level and length in PhysioNet E and F for several initializations.
PhysioNet G:
Lastly, Fig. 8 showcases TSAP’s ability to tune the level of the spike in the Extremum anomalies while the location is randomized. Note that Extremum anomalies have no length by definition. The ECG recordings in PhysioNet contain many natural spikes. As such, validation loss is low by default. Nonetheless, TSAP successfully tunes two out of four initializations and reflects – though subtly – this difference in the validation loss. Ultimately, this again leads to a well-tuned $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$ that performs strongly on $\mathcal{D}_{\mathrm{val}}$.
Figure 7: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Mean shift anomalies. Top: Given Mean shift anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (middle), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (middle), achieving high AUROC (right). Mean shift anomalies have low val. loss by default due to the subtle nature of Mean shift anomalies.
Figure 8: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Extremum anomalies. Given Extremum anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (middle), and leading to high detection AUROC performance (right). Note how Extremum anomalies have low val. loss by default due to the natural presence of spikes in ECG data.
0.A.4.2 A.4.2   Discrete Augmentation Hyperparameter Tuning
We showcased TSAP’s ability to tune the discrete hyperparameter, anomaly type, in Fig. 4 for controlled and natural TSAD tasks. Given the direct applicability and significance of discrete hyperparameter tuning in real-world contexts, we present extended results for natural TSAD tasks by additionally discussing discrete hyperparameter tuning for MoCap B (cf. Table 1).
MoCap B:
Fig. 9 shows the discrete hyperparameter tuning for the unknown anomaly type in MoCap B. TSAP was initialized twice: first with $f_{\mathrm{aug}}$ pre-trained for injecting Frequency shift anomalies, and second with $f_{\mathrm{aug}}$ pre-trained for injecting Platform anomalies. The validation loss (top) indicates a strong alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and the unlabeled $\mathcal{D}_{\mathrm{val}}$ when TSAP is initialized with Frequency shift anomalies. This is also reflected in $f_{\mathrm{det}}$’s performance on $\mathcal{D}_{\mathrm{val}}$ (middle). Visually, we can indeed confirm that Frequency shift anomalies (bottom – red) appear to be more similar to the running pattern (bottom – black) as opposed to Platform anomalies (bottom – purple).
0.A.4.3 A.4.3   Additional Ablation Studies
Figure 9: (best in color) Tuning discrete hyperparameter (anomaly type) with TSAP. For Run anomalies in MoCap B with unknown type (black), val. loss favors Frequency shift (red) that leads to high AUROC (middle) and mimics well true anomaly (bottom), and effectively rejects type Platform (purple) with poor performance.
Following Sec. 4.4, we present additional ablation studies for the second-order optimization within TSAP and the normalization of the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ obtained through $f_{\mathrm{det}}^{\mathrm{enc}}$.
Second-order Opt. Ablation:
Fig. 10 (top) shows the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when the second-order optimization is disabled. Note how the estimation process becomes highly unstable when second-order optimization is disabled. In turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ suffers severely.
Embedding Normalization Ablation:
We show the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when embedding normalization is disabled in Fig. 10 (bottom). While $a$ initialized at $0.01$ eventually leads to the correct level, the estimation process is much more volatile compared to when normalization is enabled as shown in Fig. 6 (top).
Figure 10: (best in color) Overview of additional ablation studies. Top: We disable second-order optimization in TSAP, leading to a highly unstable estimation process of $a$ (left) and poor performance of $f_{\mathrm{det}}$ (right). Bottom: We disable normalization of the embeddings in TSAP. Estimation of $a$ (left) is volatile and does not converge well, in turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ is poor in most cases (right).

### v2

@@
-
+TSA on AutoPilot: Self-tuning Self-supervised Time Series Anomaly Detection
 Abstract
@@
 A two-fold challenge for TSAD is a versatile and unsupervised model that can detect various different types of time series anomalies (spikes, discontinuities, trend shifts, etc.) without any labeled data.
-Modern neural networks have outstanding ability in modeling complex time series.
 Self-supervised models in particular tackle unsupervised TSAD by transforming the input via various augmentations to create pseudo anomalies for training.
@@
 It stands on two key components: a differentiable augmentation architecture and an unsupervised validation loss to effectively assess the alignment between augmentation type and anomaly type.
-Case studies show TSAP’s ability to effectively select the (discrete) augmentation type and associated (continuous) hyperparameters.
-In turn, it outperforms established baselines, including SOTA self-supervised models, on diverse TSAD tasks exhibiting different anomaly types.
+Case studies show TSAP’s ability to select the augmentation type and associated hyperparameters.
 1 Introduction
@@
 Thus, there exists a large body of work on time series AD as presented in various surveys [1, 2].
-Recent progress on self-supervised learning (SSL) has changed the field of AD, offering significant improvements over traditional unsupervised (or one-class) learning approaches.
-The main advantage of SSL lies in its ability to self-generate labeled samples, i.e., pseudo anomalies, within the input space.
-This enables a more focused exploration of a plausible subspace based on the semantics reflected in the pseudo anomalies, rather than an exhaustive, impractical search of the entire space.
-At the heart of SSL-based AD are data augmentation functions.
-These functions are used to create pseudo labels for the (self-)supervised training of an anomaly detector, such as by predicting whether the input is augmented or not [20], which augmentation function is used [14], or using contrastive learning [31].
-In all these approaches, the success of SSL-based AD highly depends on the degree to which the augmented data mimics the true anomalies, as outlined in [35].
-There exist few approaches for tuning data augmentation functions without labels, however, they have limitations.
-Some rely on non-differentiable validation losses [36], which are unsuitable for end-to-end learning frameworks.
-Others only address continuous hyperparameters, neglecting discrete ones, which are left for manual adjustment [34].
-For instance, in the image domain, the CutOut augmentation cannot mimic semantic class anomalies (e.g.
-cats vs.
-cars), no matter how one tunes its (continuous) hyperparameters [34], mainly due to the mismatch in the discrete choice (i.e., the augmentation type).
-Moreover, none of these efforts addresses anomaly detection for time series data, which is the focus of our work.
-In this paper we introduce TSAP, a novel approach for SSL-based time series anomaly detection (TSAD) “on autoPilot” equipped with end-to-end hyperparameter tuning.
-Effectively tuning both discrete and continuous hyperparameters of augmentation enables TSAP to be an automated anomaly detector that is most suitable for a given task.
-TSAP stands on two main components: ($i$) a differentiable parameterized augmentation model and ($ii$) an unsupervised validation loss that measures the alignment between the augmented data and the unlabeled test data, quantifying the extent to which the former mimics the true anomalies.
+Recent advances in self-supervised learning (SSL) have significantly impacted AD by surpassing traditional unsupervised (or one-class) learning approaches.
+SSL’s key advantage lies in its ability to self-generate labeled samples, or pseudo anomalies, enabling a more focused exploration of a plausible subspace rather than an exhaustive, impractical search of the entire space.
+Central to SSL-based AD are data augmentation functions, which create such pseudo labels.
+These are then used for the (self-)supervised training of an anomaly detector, such as by predicting whether the input is augmented or not [3], which augmentation function is used [4], or through contrastive learning [5].
+The success of these approaches highly depends on how well the augmented data mimics the true anomalies [6].
+In this paper we introduce TSAP, a novel approach for SSL-based time series anomaly detection (TSAD) “on autoPilot” equipped with end-to-end augmentation hyperparameter tuning.
+Tuning both discrete and continuous hyperparameters of augmentation enables TSAP to be an anomaly detector that is most suitable for a given task.
 We summarize our main contributions as follows:
   1. 1.
-    Problem: Our work is the first attempt to tune both discrete and continuous hyperparameters of data augmentation in SSL-based TSAD, without any labels at training time.
+    Problem: Our work is the first attempt to tune both discrete and continuous hyperparameters of data augmentation in SSL-based TSAD, without labels at training time.
   2. 2.
-    New TSAD Method: We propose TSAP  [^1] The code will be published once the paper is accepted., which accommodates a list of various time series anomaly types and enables automatic tuning of related hyperparameters (magnitude, duration, etc.) by a differentiable validation loss that quantifies alignment of augmented data and unlabeled test data.
+    New TSAD Method: We propose TSAP  [^1] Code and datasets available at: https://github.com/B-Deforce/TSA-on-autoPilot., which accommodates various time series anomaly types and enables automatic tuning of related hyperparameters (e.g., magnitude, duration) with a differentiable validation loss quantifying alignment between augmented and unlabeled test data.
   3. 3.
-    Effectiveness: By carefully selecting augmentation type and its (continuous) hyperparameters, our self-tuning TSAP outperforms existing unsupervised and self-supervised approaches, including the SOTA NeuTraL-AD [26] which also employs learnable augmentations, across diverse TSAD tasks.
-2 Preliminaries
-2.1 Time Series Anomaly Detection
-This work focuses on time series anomaly detection.
-Consider a univariate time series $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{K}\}$, where $\mathbf{x}$ is a sequentially ordered collection of $K$ data points.
-Each $x_{k}\in\mathbb{R}$ corresponds to a scalar observation at each time step.
-Then, let $\mathcal{D}_{\mathrm{trn}}$ be a set of training data consisting of only normal time series, and $\mathcal{D}_{\mathrm{test}}$ be a set of unlabeled test data containing both normal and anomalous time series.
-Given $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{test}}$, the problem is to detect all anomalous time series in $\mathcal{D}_{\mathrm{test}}$, i.e., to assign the accurate label $y\in\{-1,+1\}$ for each $\mathbf{x}\in\mathcal{D}_{\mathrm{test}}$.
-2.2 Self-supervised Anomaly Detectors
-Let $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ be a data augmentation function, such as rotation or cut-out in the image domain (a discrete choice), that outputs pseudo anomalies.
-A given $f_{\mathrm{aug}}$ also exhibits continuous hyperparameter(s), such as angle (for rotation) or width and height (for cutout).
-Then, a popular approach [20] is to train a detector model $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta})$ to classify between normal data $\mathcal{D}_{\mathrm{trn}}$ and augmented data $\mathcal{D}_{\mathrm{aug}}=\{f_{\mathrm{aug}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$.
-The working assumption is for $f_{\mathrm{det}}$ to predict the (unknown) anomalies in test data as augmented.
-As such, performance relies on how well the choice of augmentation type (discrete) as well as the choice(s) of its (continuous) hyperparameter(s) mimic the anomalies in $\mathcal{D}_{\mathrm{test}}$.
-There are various other ways of using SSL for anomaly detection, based on the specific objective function and how the pseudo labels are generated [14, 31], which are similar to each other in principle.
-2.3 Data Augmentation on Time Series
-To apply SSL to time series data, we need an augmentation function specifically designed for time series.
-For a given time series $\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}$, $f_{\mathrm{aug}}(\mathbf{x};\mathbf{a})$ transforms $\mathbf{x}$ based on hyperparameters $\mathbf{a}\in\mathcal{A}^{P}$, $P\geq 1$, where $\mathcal{A}^{P}$ represents the domain of all possible hyperparameter values.
-For time series data, augmentations could involve mean shifts, trend injections, or other transformations relevant to the time-series domain, with $\mathcal{A}^{P}$ encompassing the respective hyperparameter space for a given transformation.
-We describe augmentation in more detail in Sec.
-3.1.
-2.4 Wasserstein Distance
-The Wasserstein distance [1], a distance measure between probability distributions, of order $p$ between any two marginal distributions $\mu$ and $\rho$ is given as
-$W_{p}(\mu,\nu)=\biggl{(}\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}[d(x,y)^{p}]\biggr{)}^{1/p},$ (1)
-[…]
+    Effectiveness: By carefully selecting augmentation type and its (continuous) hyperparameters, TSAP outperforms existing unsupervised and self-supervised approaches, including the SOTA NeuTraL-AD [7] which also employs learnable augmentations.
 2 TSAP: Time Series Anomalies on AutoPilot
@@
 Left: Offline trained, differentiable augmentation model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ takes as input the normal data and augmentation hyperparameter(s) $\mathbf{a}$, and outputs pseudo-anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$.
-Right: Self-tuning engine incorporates the pre-trained $f_{\mathrm{aug}}$ (with parameters $\boldsymbol{\phi}$ frozen), alternating between two phases: ($i$) detection phase – given $\mathbf{a}^{(t)}$ at iteration $t$, estimate parameters $\boldsymbol{\theta}^{(t)}$ of detector $f_{\mathrm{det}}$ (consisting of Encoderθ and discriminator MLPθ), by optimizing $\mathcal{L}_{\mathrm{trn}}$ (classification loss); ($ii$) alignment phase – given $f_{\mathrm{det}}^{\mathrm{enc}}(\cdot;\boldsymbol{\theta}^{(t)})$, update augmentation (governed by $\mathbf{a}$) to better align the embedded time series $\mathbf{z}_{\mathrm{trn}}\cup\mathbf{z}_{\mathrm{aug}}$ with $\mathbf{z}_{\mathrm{val}}$ in the learned discriminative space.
-Note that $\mathbf{x}_{\mathrm{val}}$ contains both normal and anomalous time series, but labels are not known or used at any point during training time.
+Right: Self-tuning engine incorporates the pre-trained $f_{\mathrm{aug}}$ (with parameters $\boldsymbol{\phi}$ frozen), alternating between two phases as outlined in Algo.
+1.
 There are two notable challenges that need to be addressed for automatic selection of both discrete and continuous hyperparameters for SSL-based TSAD:
-Challenges
-  • C1.
+  1. C1.
     Differentiable Augmentation: Developing an augmentation function that is differentiable with respect to its hyperparameters, enabling gradient-based optimization.
-  • C2.
-    Comparable Validation Loss: Formulating a validation loss that effectively quantifies alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{test}}$ while being comparable across different hyperparameter initializations.
-To address C1 and C2, we center our framework around two main ideas:
-Main ideas
-  • I1.
-    Differentiable Augmentation Module: TSAP implements augmentation function as an Encoder-Decoder neural network, $f_{\mathrm{aug}}$ parameterized by $\boldsymbol{\phi}$, capable of approximating the anomaly-generating mechanism conditioned on $\mathbf{a}\in\mathcal{A}^{P}$.
-Importantly, this module is pre-trained independently and prior to the initiation of I2, establishing it as an offline component of the framework.
-  • I2.
-    Self-Tuning Module: At test time online, TSAP iteratively refines the detector $f_{\mathrm{det}}$’s parameters $\boldsymbol{\theta}$ as well as augmentation hyperparameters $\mathbf{a}$, through alternating detection and alignment phases.
+  2. C2.
+    Comparable Validation Loss: Formulating a validation loss that quantifies alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{test}}$ while being comparable across different hyperparameter initializations.
+To address C1 and C2, we center our framework around two main ideas.
+First, TSAP employs a differentiable augmentation module implemented as an Encoder-Decoder neural network, $f_{\mathrm{aug}}$, parameterized by $\boldsymbol{\phi}$.
+This module approximates the anomaly-generating mechanism, conditioned on $\mathbf{a}\in\mathcal{A}^{P}$, $P\geq 1$, where $\mathcal{A}^{P}$ represents the domain of all possible hyperparameter values (e.g., magnitude, duration, …).
+Importantly, this module is pre-trained independently, establishing it as an offline component of the framework.
+Second, at test time online, TSAP iteratively refines the detector $f_{\mathrm{det}}$’s parameters $\boldsymbol{\theta}$ as well as augmentation hyperparameters $\mathbf{a}$, through alternating detection and alignment phases.
 Alignment is performed on part of the unlabeled $\mathcal{D}_{\mathrm{test}}$, referred to as $\mathcal{D}_{\mathrm{val}}$.
-Based on the key ideas I1 and I2, we propose TSAP, a self-tuning self-supervised TSAD framework illustrated in Fig. 1.
-Figure 2: Examples of six different types of time series anomalies; (black) original real-world time series, (red) pseudo anomalies generated by $g$.
-The goal for $f_{\mathrm{aug}}$ is to learn to inject anomalies into inliers, conditional on the discrete (anomaly type) as well as continuous hyperparameters such as level (magnitude), location (start point) and length (see bottom right).
-More details are presented in Sec.
-3.1 and Appx.
-0.A.1.
+Based on this, we propose TSAP, a self-tuning self-supervised TSAD framework illustrated in Fig.
+1.
 2.1 Differentiable Augmentation Module
@@
 We carefully consider accommodating six types of anomalies that are common in real-world time series; namely, trend, extremum, amplitude, mean shift, frequency shift, and platform.
-We provide a detailed description of the anomalies in Appx.
-0.A.1 and for brevity, illustrate them by examples in Fig.
-2 with red lines.
-Each type of anomaly has three hyperparameters; including its starting position (location), duration (length), and severity (level) as illustrated in Fig.
-2 (bottom right).
+We provide a detailed description with visual examples of the anomalies in Appx.
+A.
+Each type of anomaly has three hyperparameters; including its starting position (location), duration (length), and severity (level).
 Extremum captures a spike and only has two hyperparameters as its duration is always $1$.
-Based on $\mathcal{D}_{\mathrm{trn}}$, the anomaly generation scheme $g$ creates an augmented dataset $\mathcal{D}_{\mathrm{aug}}=\{g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})\mid\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}},\mathbf{a}\sim\mathcal{A}^{P}\}$, where $\mathbf{a}$ is the vector of augmentation hyperparameters that are uniformly randomly sampled from the domain $\mathcal{A}^{P}$.
+Based on $\mathcal{D}_{\mathrm{trn}}$, the anomaly generation scheme $g$ creates an augmented dataset $\mathcal{D}_{\mathrm{aug}}=\{g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})\mid\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}},\mathbf{a}\sim\mathcal{A}^{P}\}$, where $\mathbf{a}$ is the vector of augmentation hyperparameters that are uniformly randomly sampled from the hyperparameter domain $\mathcal{A}^{P}$.
 Model Design
-To build an augmentation model $f_{\mathrm{aug}}$ for time series, first we use a convolutional neural network (CNN) to encode or embed the input time series $\mathbf{x}_{\mathrm{trn}}$ into the feature map $\mathbf{z}_{\mathrm{trn}}$.
+To build an augmentation model $f_{\mathrm{aug}}$ for time series, we use a convolutional neural network (CNN) to encode the input time series $\mathbf{x}_{\mathrm{trn}}$ into the feature map $\mathbf{z}_{\mathrm{trn}}$.
 We then encode the augmentation hyperparameters $\mathbf{a}$ into $\mathbf{z}_{\mathrm{a}}$, which has the same shape as $\mathbf{z}_{\mathrm{trn}}$, with a multilayer perceptron (MLP).
 Since the feature map $\mathbf{z}_{\mathrm{trn}}$ generated by the CNN encoder keeps the positional information of the original time series, adding $\mathbf{z}_{\mathrm{a}}$ to $\mathbf{z}_{\mathrm{trn}}$ ensures that only the part with the desired location and length in $\mathbf{z}_{\mathrm{aug}}$ is manipulated.
-To ensure that the feature maps $\mathbf{z}_{\mathrm{aug}}$ and $\mathbf{z}_{\mathrm{trn}}$ are in the same embedding space, they share the same decoder to reconstruct back to the time series $\mathbf{x}_{\mathrm{aug}}=g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ and $\mathbf{x}_{\mathrm{trn}}$, respectively.
+To ensure that the feature maps $\mathbf{z}_{\mathrm{aug}}$ and $\mathbf{z}_{\mathrm{trn}}$ are in the same embedding space, they share the same decoder to reconstruct back to the time series $\mathbf{x}_{\mathrm{aug}}=g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ and $\mathbf{x}_{\mathrm{trn}}$.
 As such, the loss function of $f_{\mathrm{aug}}$ is based on the reconstruction of both $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{aug}}$:
 $\mathcal{L}_{\mathrm{aug}}=\sum_{\begin{subarray}{c}\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}}\\ \mathbf{a}\sim\mathcal{A}^{P}\end{subarray}}\big{(}{(\mathbf{x}_{\mathrm{trn}}-\mathbf{\tilde{x}}_{\mathrm{trn}})}^{2}+{(g(\mathbf{x}_{\mathrm{trn}},\mathbf{a})-\mathbf{\tilde{x}}_{\mathrm{aug}})}^{2}\big{)}$ (1)
-where $\mathbf{\tilde{x}}_{\mathrm{aug}}$ denotes the output of $f_{\mathrm{aug}}(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ for the same value of the hyperparameters sampled by $g$ (see Fig. 1 left).
+where $\mathbf{\tilde{x}}_{\mathrm{aug}}$ denotes the output of $f_{\mathrm{aug}}(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ for the same $\mathbf{a}$ as sampled by $g$ (see Fig. 1 left).
+2.2 Self-Tuning Module
 Algorithm 1 Self-Tuning Module of TSAP
@@
 14:Return $\mathbf{a}^{*}$ and $\boldsymbol{\theta}^{*}$
-3.2 Self-Tuning Module
 Central to TSAP is the self-tuning module, which operates by iteratively refining the detector’s parameters, $\boldsymbol{\theta}$, and the augmentation hyperparameters, $\mathbf{a}$.
@@
 1.
-Phase ($i$): Detection
-This phase focuses on estimating the parameters $\boldsymbol{\theta}^{(t)}$ of the detector $f_{\mathrm{det}}$ (comprising of an encoder $f_{\mathrm{det}}^{\mathrm{enc}}$ and a discriminator $f_{\mathrm{det}}^{\mathrm{mlp}}$) by minimizing the cross-entropy loss $\mathcal{L}_{\mathrm{trn}}$.
-This aims to classify between the normal samples $\mathbf{x}_{\mathrm{trn}}$ and the augmented pseudo anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$ by their embeddings $\mathcal{Z}_{\mathrm{trn}}$ and $\mathcal{Z}_{\mathrm{aug}}$, where $\mathcal{Z}_{\mathrm{trn}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ and $\mathcal{Z}_{\mathrm{aug}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x};\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ denote the embeddings of the training data and augmented data, respectively, given the current fixed $\mathbf{a}^{(t)}$ at iteration $t$.
-Note that the parameters $\boldsymbol{\phi}$ of the augmentation model $f_{\mathrm{aug}}$ are frozen throughout this phase.
-Phase ($ii$): Alignment
+Phase ($i$) Detection
+This phase is focused on estimating the parameters $\boldsymbol{\theta}^{(t)}$ of the detector $f_{\mathrm{det}}$ (comprising of an encoder $f_{\mathrm{det}}^{\mathrm{enc}}$ and a discriminator $f_{\mathrm{det}}^{\mathrm{mlp}}$) by minimizing the cross-entropy loss $\mathcal{L}_{\mathrm{trn}}$.
+This aims to classify between the normal samples $\mathbf{x}_{\mathrm{trn}}$ and the augmented pseudo anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$ by their embeddings $\mathcal{Z}_{\mathrm{trn}}$ and $\mathcal{Z}_{\mathrm{aug}}$, where $\mathcal{Z}_{\mathrm{trn}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ and $\mathcal{Z}_{\mathrm{aug}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x};\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ denote the embeddings of the training data and augmented data, given the current fixed $\mathbf{a}^{(t)}$ at iteration $t$.
+Note that the parameters $\boldsymbol{\phi}$ of $f_{\mathrm{aug}}$ are frozen throughout this phase.
+Phase ($ii$) Alignment
 Subsequently, the alignment phase adjusts $\mathbf{a}$ to optimize the unsupervised differentiable validation loss $\mathcal{L}_{\mathrm{val}}$, computed based on the embeddings from the now-updated $f_{\mathrm{det}}^{\mathrm{enc}}$.
 $\mathcal{L}_{\mathrm{val}}$’s objective is to measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space, as expressed by the Wasserstein distance in Eq.
-(1).
-This metric is chosen for its effectiveness in capturing the overall distributional discrepancies between datasets, offering a more complete comparison than mere point-wise metrics [1], which is especially important in TSAD given the often subtle nature of time series anomalies.
-Note that the embeddings are normalized to ensure scale invariance before being passed to $\mathcal{L}_{\mathrm{val}}$, so as to avoid the trivial solution of achieving alignment by setting all embeddings to the zero vector [34].
-Lastly, as the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ in $\mathcal{L}_{\mathrm{val}}$ are obtained through the updated $f_{\mathrm{det}}^{\mathrm{enc}}$, the optimization process needs to track the change in $\boldsymbol{\theta}$ caused by the update of $\mathbf{a}$.
+(2), Appx.
+B.1.
+Note that the embeddings are normalized to ensure scale invariance before being passed to $\mathcal{L}_{\mathrm{val}}$.
+This avoids the trivial solution of achieving alignment by setting all embeddings to $\mathbf{0}$ [8].
+As the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ in $\mathcal{L}_{\mathrm{val}}$ are obtained through the updated $f_{\mathrm{det}}^{\mathrm{enc}}$, the optimization needs to track the change in $\boldsymbol{\theta}$ caused by the update of $\mathbf{a}$.
 As such, TSAP employs a second-order optimization process, similar to [8].
@@
 1 describes continuous hyperparameter tuning, the discrete hyperparameter (i.e.
-augmentation type) selection, corresponding to a specific anomaly type, is done through grid search as the number of anomaly types is finite.
+augmentation/anomaly type) selection is done through grid search as the number of anomaly types is finite.
 Hence, we initialize TSAP for different augmentation types and compare $\mathcal{L}_{\mathrm{val}}$ across types to select the one that yields the best alignment.
@@
 3 Experiments
-In this section, we aim to answer the following questions:
-  1. Q1.
-    Quantitative Comparison: How well does TSAP perform on TSAD against established baselines?
-  2. Q2.
-    Qualitative Analysis: How well does TSAP tune the augmentation hyperparameters?
-  3. Q3.
-    Ablation Studies: Which design decisions in TSAP drive its effectiveness?
-4.1 Setup
-Datasets
-We evaluate TSAP on six distinct TSAD tasks.
-Four of these are conducted in a controlled environment, while the remaining two are natural.
-In the controlled setting, the anomaly types are manually injected in $\mathcal{D}_{\mathrm{test}}$ based on the types discussed in Sec.
-3.1.
-The controlled environment allows for a thorough assessment of TSAP’s ability to perform continuous hyperparameter tuning.
-For the natural environment, the anomaly types in $\mathcal{D}_{\mathrm{test}}$ are a priori unknown and it is the goal of TSAP to find the anomaly type that yields best alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ (part of $\mathcal{D}_{\mathrm{test}}$), expressed by $\mathcal{L}_{\mathrm{val}}$.
-The details of the datasets and anomaly types are provided in Appx.
-0.A.1 and 0.A.2.
-For controlled tasks, we use the 2017 PhysioNet Challenge dataset [9], which comprises real-world ECG recordings.
-Table 1 shows a selection of four different controlled TSAD tasks, constructed by manually injecting different anomaly profiles into the PhysioNet dataset.
-For PhysioNet A and B, given the anomaly type (Platform), the task is to infer or tune, respectively, the hyperparameter(s) level only and both level and length, while anomaly location is random (hence not tuned).
-For PhysioNet C and D, the respective tuning tasks are the same but for a different anomaly type (Trend).
-We choose Platform and Trend anomalies to represent, respectively, range-bound and range-exceeding anomalies.
-Additional variations of PhysioNet with other injected anomaly types, such as Mean shift and Extremum, are discussed in Appx.
-0.A.2.
-The natural TSAD tasks are derived from the CMU Motion Capture (MoCap) dataset [^2] http://mocap.cs.cmu.edu/.
-We consider the walking signal as normal data, and the signals of jumping and running as anomalies.
-To generate normal signals, we stitch the walking signals by identifying the start and end points of each gait phase and add random noise; whereas to generate anomalous ones, we stitch walking or running signals at a random location in the normal signal.
-This yields two distinct TSAD tasks as shown in Table 1.
-Different from PhysioNet A–D where we only tune the continuous hyperparameter(s) for the given (discrete) anomaly type, here for MoCap A and B, we aim to tune both the unknown anomaly type that corresponds to Jump and Run behavior, respectively, as well as the (continuous) hyperparameter level while location and length take random values.
-Table 1: Anomaly profile of different TSAD tasks.
-Profile PhysioNet A PhysioNet B PhysioNet C PhysioNet D MoCap A MoCap B
-Type Platform Platform Trend Trend Jump Run
-Level Fixed Fixed Fixed Fixed Fixed Fixed
-Location Random Random Random Random Random Random
-Length Random Fixed Random Fixed Random Random
-Baselines
-We compare TSAP with a selection of established baselines, including both traditional and deep learning-based methods that have demonstrated efficacy in TSAD [29].
-Among the traditional methods, we utilize three with different modeling approaches; namely, One-Class Support Vector Machines (OC-SVM) [30]; Local Outlier Factor (LOF) [7]; and Isolation Forest (IF) [21].
-On the deep learning side, we benchmark against the Encoder-Decoder LSTM (EncDec-LSTM) [24]; the Spectral Residual Convolutional Neural Network (SR-CNN) [28]; and the Unsupervised Anomaly Detection (USAD) for TSAD [2].
-Lastly, we include a state-of-the-art competing method which learns augmentations in the embedding space, called Neural Transformation Learning for (TS)AD (NeuTraL-AD) [26].
-This diverse set of baselines allows for a comprehensive analysis across different approaches within the TSAD domain.
-Model Configurations
-The Encoderϕ in $f_{\mathrm{aug}}$ and Encoderθ in $f_{\mathrm{det}}$ are constructed using 1D CNN blocks [4] (transposed 1D CNN for Decoderϕ) for efficient temporal feature extraction.
-We carefully choose the number of epochs $T$ to allow sufficient time for the convergence of $\mathbf{a}$, with empirical evidence suggesting that $T=100$ typically suffices.
-For the number of inner-loops $L$, we opt for $L=5$, aligned with [34], such that $f_{\mathrm{det}}$ has adequate time to learn effective discriminative embeddings for $\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{trn}}$.
-A detailed overview of the configuration for TSAP and all baselines is provided in Appx.
-0.A.3.
-Evaluation Metrics
-Our method calculates anomaly scores on an entire sequence level $\mathbf{x}$, similar to [26].
-This is a different set-up compared to novelty detection in time series which typically operates on a point level.
-Detection on a sequence level can be especially important to spot long-term patterns (e.g.
-Trend anomalies).
-As such, we use the Area Under the Receiver Operating Characteristic Curve (AUROC) and the $F_{1}$ score as key performance metrics to quantify detection capability of anomalous sequences.
-All results are reported on the unseen $\mathcal{D}_{\mathrm{test}}$.
-In addition, we determine the optimal $F_{1}$ score, by enumerating all possible thresholds, given by the anomaly scores for a given segment $\mathbf{x}$.
-We then compute the corresponding precision and recall for each threshold and select those that yield the highest $F_{1}$.
-As such, AUROC provides a balanced view whereas $F_{1}$ shows the optimal scenario.
-Both metrics range from 0 to 1, with higher values indicating superior performance.
-4.2 Quantitative Results
 Table 1: Detection peformance of baselines w.r.t.
-$F_{1}$ and AUROC on test data across six TSAD tasks.
-Four tasks are performed in a controlled setting (anomaly type is manually injected and corresponds to anomaly types described in Sec.
-3.1), based on the PhysioNet ECG data.
-Remaining two TSAD tasks exhibit natural anomalies (real-world anomaly type which is unknown), based on the MoCap data.
-TSAP outperforms most baselines and has the best (lowest) average rank w.r.t.
-both $F_{1}$ and AUROC, with low standard deviation (in parentheses).
+$F_{1}$ and AUROC on test data across six TSAD tasks (cf.
+Appx.
+C.1, Table 3).
+TSAP has the lowest average rank w.r.t.
+$F_{1}$ and AUROC, with low std.
+dev.
 Methods PhysioNet A PhysioNet B PhysioNet C PhysioNet D MoCap A MoCap B Avg. Rank
 $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC
-OC-SVM 0.182 0.468 0.182 0.472 0.373 0.803 0.393 0.806 1.000 1.000 0.546 0.806 5.7 (2.6) 5.3 (2.3)
-LOF 0.999 1.000 0.999 1.000 0.354 0.738 0.358 0.725 0.196 0.506 0.221 0.577 5.5 (2.7) 5.2 (3.3)
-IF 0.255 0.587 0.232 0.576 0.183 0.402 0.182 0.356 0.864 0.965 0.342 0.758 6.2 (1.6) 6.3 (1.4)
-EncDec-LSTM 0.190 0.508 0.190 0.508 0.415 0.812 0.442 0.819 0.980 0.999 0.909 0.996 4.8 (1.0) 5.00 (0.9)
-SR-CNN 0.965 0.990 0.964 0.998 0.983 0.999 0.991 0.999 0.302 0.700 0.214 0.512 3.8 (3.0) 4.0 (2.8)
-USAD 0.183 0.425 0.184 0.428 0.430 0.822 0.409 0.828 1.000 1.000 1.000 1.000 4.2 (2.7) 4.3 (3.1)
-NeuTraL-AD 0.211 0.732 0.263 0.679 0.561 0.868 0.526 0.862 1.000 1.000 1.000 1.000 2.8 (1.6) 2.7 (1.4)
-TSAP (ours) 1.000 1.000 1.000 1.000 0.973 0.999 0.991 0.998 0.889 0.969 1.000 1.000 1.8 (1.6) 1.8 (1.6)
-Table 2 provides the detection results for all six TSAD tasks described in Sec.
-4.1.
-TSAP ranks the best overall in terms of both $F_{1}$ and AUROC.
-This shows that detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$, trained through the alternating mechanism of TSAP, is able to generalize to unseen and unlabeled anomalies in $\mathcal{D}_{\mathrm{test}}$.
-While some competing methods perform strongly on a subset of tasks, they lack consistency across all TSAD tasks.
-Among the traditional baselines, which perform subpar as compared to deep TSAD approaches, LOF performs very well in PhysioNet A and B, relatively poorly on C and D, but exhibits near-random performance on MoCap.
-This discrepancy can be attributed to the nature of the anomalies in the PhysioNet data, which are manually injected and often characterized by pronounced and abrupt changes as shown in Fig.
-2.
-These abrupt changes lead to large differences in local density, which makes LOF thrive.
-In contrast, the MoCap dataset contains anomalies that emerge more subtly, as one gradually transitions from walking to running or jumping, reflecting more natural variations in the data which, in turn, lead to more subtle differences in local density.
-Similarly, SR-CNN performs strongly on the PhysioNet TSAD tasks, yet its effectiveness also diminishes drastically on MoCap tasks.
-SR-CNN exploits the abrupt changes present in the PhysioNet data but fails to detect the more subtle anomalies in MoCap.
-This suggests that, while SR-CNN and LOF work well for certain types of anomalies, their utility may be limited in scenarios where anomalies are more subtle.
-On the contrary, reconstruction-based baselines such as EncDec-LSTM and USAD face challenges on PhysioNet due to its high variability among inliers, which complicates the task of accurate data reconstruction.
-Yet, these methods excel with the MoCap dataset, where its consistent near-periodic pattern (Fig.
-2, bottom) lends itself to more reliable reconstruction, thereby enhancing anomaly detection performance.
-Similarly, NeuTraL-AD, a state-of-the-art augmentation-based baseline, demonstrates proficiency in properly augmenting the inlier data within MoCap.
-However, its performance on the PhysioNet variations reveals some weaknesses.
-NeuTraL-AD struggles with the high variability in the PhysioNet inliers, inherent to real-world ECG signals.
-This suggests that the augmentation functions considered in NeuTraL-AD lack robustness when inliers are inherently noisy.
-We remark that only TSAP provides robust and consistent performance across all TSAD tasks, showcasing the effectiveness and generalizability of our proposed method.
-4.3 Qualitative Results
-Figure 3: (best in color) Tuning continuous augmentation hyperparameter(s) with TSAP.
-Top: Given Platform anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val.
-loss (center), and leading to high detection performance (right).
-Bottom: Multiple continuous hyperparameters, here both level and length, are accurately tuned to near true values (left), as guided by minimizing the val.
-loss (center), achieving high AUROC (right).
-Notice that the diverged results in both cases associate with high (or hard-to-optimize) val.
-loss, which help us effectively reject low performance models.
-A key contributor to TSAP’s consistent performance is the validation loss through which TSAP automatically learns the augmentation hyperparameters $\mathbf{a}$.
-Once $\mathbf{a}$ is determined, the task essentially reduces to a supervised learning problem.
-Next, we show that TSAP not only effectively tunes the continuous augmentation hyperparameters $\mathbf{a}$, but also that its validation loss guides the accurate selection of the discrete hyperparameter (i.e.
+OC-SVM 0.182 0.468 0.182 0.472 0.373 0.803 0.393 0.806 1.000 1.000 0.546 0.806 7.2 (3.7) 6.8 (3.3)
+LOF 0.999 1.000 0.999 1.000 0.354 0.738 0.358 0.725 0.196 0.506 0.221 0.577 6.8 (3.9) 6.5 (4.4)
+ARIMA 0.885 0.960 0.829 0.965 0.991 0.999 0.999 0.999 0.870 0.955 0.225 0.537 4.3 (3.2) 4.7 (3.7)
+IF 0.255 0.587 0.232 0.576 0.183 0.402 0.182 0.356 0.864 0.965 0.342 0.758 8.8 (1.7) 8.7 (1.9)
+MP 0.812 0.743 0.812 0.744 0.280 0.712 0.284 0.734 1.000 1.000 1.000 1.000 5.0 (3.6) 4.8 (3.4)
+EncDec-LSTM 0.190 0.508 0.190 0.508 0.415 0.812 0.442 0.819 0.980 0.999 0.909 0.996 6.5 (2.0) 6.7 (1.9)
+SR-CNN 0.965 0.990 0.964 0.998 0.983 0.999 0.991 0.999 0.302 0.700 0.214 0.512 5.2 (4.2) 4.8 (4.5)
+USAD 0.183 0.425 0.184 0.428 0.430 0.822 0.409 0.828 1.000 1.000 1.000 1.000 5.5 (4.0) 5.7 (4.5)
+NeuTraL-AD 0.211 0.732 0.263 0.679 0.561 0.868 0.526 0.862 1.000 1.000 1.000 1.000 4.2 (2.9) 3.8 (2.5)
+TimeGPT 0.327 0.714 0.318 0.711 0.218 0.580 0.217 0.525 0.348 0.743 0.385 0.683 8.0 (1.9) 8.3 (1.6)
+TSAP (ours) 1.000 1.000 1.000 1.000 0.973 0.999 0.991 0.998 0.889 0.969 1.000 1.000 2.3 (2.0) 2.2 (2.0)
+We evaluate TSAP on six distinct TSAD tasks.
+Four of these are conducted in a controlled environment, using the 2017 ECG PhysioNet Challenge dataset [9].
+The remaining two are natural anomalies, derived from the CMU Motion Capture (MoCap) dataset.
+In the former, the anomaly types are manually injected in $\mathcal{D}_{\mathrm{test}}$ based on the types discussed in Sec.
+2.1.
+For the latter, the anomaly types in $\mathcal{D}_{\mathrm{test}}$ are unknown and it is the goal of TSAP to find the type that yields best alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ (part of $\mathcal{D}_{\mathrm{test}}$), expressed by $\mathcal{L}_{\mathrm{val}}$.
+See Appx.
+C for further details.
+Quantitative Results
+Table 1 shows the detection results for these six TSAD tasks.
+TSAP ranks the best overall in both $F_{1}$ and AUROC, showing that detector $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$, trained through the alternating mechanism of TSAP, is able to generalize to unseen, unlabeled anomalies in $\mathcal{D}_{\mathrm{test}}$.
+While competing methods perform well on certain tasks, they lack consistency across all TSAD tasks.
+LOF, for instance, thrives on PhysioNet A and B due to its abrupt, manually injected anomalies, but fails on MoCap, where anomalies emerge more gradually as activities transition naturally.
+SR-CNN and ARIMA show similar behavior.
+See Appx.
+B.2B.3 for model/baseline configuration details.
+Reconstruction-based methods like EncDec-LSTM and USAD struggle on PhysioNet due to high variability among inliers, yet excel on MoCap, which exhibits consistent, near-periodic patterns.
+NeuTraL-AD, despite its augmentation-based approach, struggles with PhysioNet’s real-world ECG signal variability, suggesting its augmentation functions lack robustness in noisy conditions.
+We remark that only TSAP provides robust and consistent performance across all TSAD tasks.
+Qualitative Results
+Figure 2: Tuning continuous augmentation hyperparameter(s) with TSAP.
+Top: For Platform anomalies at true level (red dashed line), various initializations converge near the true value (left) by minimizing val.
+loss (center), which leads to high detection performance (right).
+Bottom: Multiple continuous hyperparameters (level and length) converge near true values (left) through minimizing val.
+loss (center), resulting in high AUROC (right).
+Diverged results correspond to high val.
+loss, aiding in rejecting low-performance models.
+Figure 3: Tuning the discrete hyperparameter (anomaly type) with TSAP.
+Top: For true Trend anomaly (black), val.
+loss favors Trend (green) and Mean shift (blue), both achieving high AUROC (middle) and strong resemblance (right), while rejecting Platform (orange) with poor performance.
+Bottom: For Jump anomalies in MoCap A with unknown type (black), val.
+loss favors Platform (purple), leading to high AUROC (middle) and similarity to the true anomaly (right), while rejecting Frequency (red) with poor performance.
+Key to TSAP’s consistent performance is $\mathcal{L}_{\mathrm{val}}$ through which TSAP automatically learns the augmentation hyperparameters $\mathbf{a}$.
+Once $\mathbf{a}$ is determined, the task reduces to a supervised learning problem.
+Next, we show that TSAP not only effectively tunes the continuous augmentation hyperparameters $\mathbf{a}$, but also that $\mathcal{L}_{\mathrm{val}}$ guides the accurate selection of the discrete hyperparameter (i.e.
 anomaly type).
-Controlled Environment
 Consider PhysioNet A, where we aim to tune the continuous hyperparameter $a$, i.e.
@@
 We observe that the initialization for $a\in\{-0.4,0.6\}$ leads to the true $a$$=$$0.2$ (left).
-Simultaneously, the validation loss drops substantially once TSAP has arrived at the true $a$ (middle).
-This is also reflected in the performance of $f_{\mathrm{det}}$ on part of the unlabeled test data $\mathcal{D}_{\mathrm{val}}$ which soars upon estimation of the true $a$ (right).
-Conversely, the initialization for $a\in\{0.8,-0.8\}$ leads to a high validation loss, indicating poor alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$.
+Simultaneously, $\mathcal{L}_{\mathrm{val}}$ drops substantially once TSAP has arrived at the true $a$ (center).
+This is also reflected in the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ which soars upon estimation of the true $a$ (right).
+Conversely, the initialization for $a\in\{0.8,-0.8\}$ leads to a high $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$.
 Indeed, the performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ now suffers from poor alignment.
@@
 3 (bottom) demonstrates TSAP’s ability to accurately estimate the level and length.
-While this is a by-product of our method, there are several real-world use-cases that can directly benefit from accurately learning the anomaly profile, from industrial equipment monitoring, to network security, and healthcare monitoring.
-This demonstrates TSAP’s versatility given its capability of estimating continuous hyperparameters.
-See Appx.
-0.A.4 for additional qualitative results on other anomaly types such as Trend, Mean shift, and Spike.
 Further, Fig.
 3 (top) showcases TSAP’s ability to perform discrete hyperparameter selection.
-Here, TSAP has been initialized and trained with three different anomaly types (Mean shift, Platform, Trend) on PhysioNet C (true anomaly type is Trend).
-The validation loss clearly showcases a misalignment between the Platform and Trend types (left), which is also reflected in the AUROC of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ (middle).
-Note how the Mean shift anomaly type has a low $\mathcal{L}_{\mathrm{val}}$ towards the end of the training epochs, which is also reflected in the AUROC on $\mathcal{D}_{\mathrm{val}}$.
-Indeed, comparing TSAP tuned for Mean shift anomalies and for Trend anomalies show strong resemblances with the true underlying anomaly type (right).
-This demonstrates that the true underlying anomaly type is not necessarily the only type that yields high alignment, and in turn, a high-performing detector.
-Natural Environment
-In MoCap datasets, the anomaly types are a priori unknown.
-As such, we initialize TSAP with different augmentation types (Frequency and Platform) to perform discrete hyperparameter selection.
+Here, TSAP has been initialized and trained with three anomaly types (Mean shift, Platform, Trend) on PhysioNet C (true anomaly type: Trend).
+$\mathcal{L}_{\mathrm{val}}$ indicates a misalignment between the Platform and Trend types (left), also reflected in the AUROC of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ (center).
+Note how the Mean shift anomaly type has a low $\mathcal{L}_{\mathrm{val}}$ at the end of the training epochs, reflected in the high AUROC on $\mathcal{D}_{\mathrm{val}}$.
+This shows that the true underlying anomaly type is not necessarily the only type that yields high alignment, and in turn a high-performing detector.
+In MoCap datasets, where anomaly types are a priori unknown, TSAP is initialized with different augmentation types (Frequency and Platform) to perform discrete hyperparameter selection.
 Fig.
-4 (bottom) highlights its effectiveness as the validation loss clearly prefers one type over the other.
+3 (bottom) highlights its effectiveness as $\mathcal{L}_{\mathrm{val}}$ clearly prefers one type over the other.
 Indeed, the natural anomalies defined by jumping signals in MoCap A have close resemblance to platform anomalies.
-Discrete hyperparameter optimization for MoCap B is similar, as given in Appx.
-0.A.4.
-Figure 4: (best in color) Tuning the discrete hyperparameter (anomaly type) with TSAP.
-Top: Given true Trend anomaly (black), val.
-loss favors both Trend (green) and Mean shift (blue) type, both with high AUROC (middle) and high resemblance (right), and effectively rejects type Platform (orange) with poor performance.
-Bottom: For Jump anomalies in MoCap A with unknown type (black), val.
-loss favors type Platform (purple) that leads to high AUROC (middle) and mimics well the true anomaly (right), and effectively rejects type Frequency (red) with poor performance.
-4.4 Ablation Studies
-We showcase two ablation studies on the controlled PhysioNet data to support various design strategies of TSAP.
-Additional ablation studies on embedding normalization and second-order optimization are provided in Appx.
-0.A.4.
-Validation Loss
-In Fig.
-5 (top), we illustrate the level estimation for PhysioNet C under the condition where our Wasserstein-based $\mathcal{L}_{\mathrm{val}}$ is substituted with a point-wise metric, as used in [34].
-This comparison shows that a point-wise validation loss tends to favor solutions where the level of the Trend anomaly approximates zero, essentially neutralizing the anomaly.
-Although this might produce high alignment, it leads to poor $f_{\mathrm{det}}$ performance in $\mathcal{D}_{\mathrm{val}}$ (right).
-This shows that the distributional characteristics captured by our $\mathcal{L}_{\mathrm{val}}$ are a key contributing factor to the success of TSAP.
-Figure 5: (best in color) Overview of ablation studies.
-Top: TSAP’s $\mathcal{L}_{\mathrm{val}}$ is replaced with a point-wise val.
-loss leading to an erroneous estimation of $a$ (left) and poor performance (right).
-Bottom: TSAP’s self-tuning module is disabled and $a$ is instead randomized.
-Val.
-loss indicates poor alignment, reflected in poor performance.
-Randomization vs. Tuning
-In Fig.
-5 (bottom), the self-tuning module is disabled for PhysioNet C, where level is instead randomized (along with location, and length).
-We observe substantially higher $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment.
-In turn, $f_{\mathrm{det}}$ struggles to detect the unlabeled anomalies in $\mathcal{D}_{\mathrm{val}}$.
-This clearly demonstrates the utility of TSAP’s systematic hyperparameter (self-) tuning over random choice.
-5 Related Work
-Self-supervised Learning (SSL) for Anomaly Detection:
-SSL has emerged as a significant approach in machine learning.
-Foundation models [6] like large language models [25, 27], whose training heavily relies on SSL, have shaken up the world with outstanding performance.
-SSL has also transformed representation learning and significantly boosted several tasks in NLP, computer vision, recommender systems and medicine [3, 19].
-SSL is especially attractive for unsupervised problems like anomaly detection (AD) [16], because of its nature to create proxy tasks and loss functions without any labels.
-There are various types of SSL-based methods on AD [22], but the core idea is to use data augmentation functions [8, 14, 20] on inliers to create pseudo anomalies, and learn a classifier that can detect the pseudo anomalies.
-Time Series Anomaly Detection:
-There exists a large body of work on TSAD, for which we refer to several surveys [5, 15].
-Recently, SSL has been widely studied as the main approach to address TSAD problems [32, 37].
-BeatGAN [43] and RobustTAD [13] use data augmentation to enrich training data.
-COUTA [33] proposes three different augmentations, respectively mimicking local, contextual and collective anomalies.
-NeuTraL-AD [26] propose several augmentation functions that are an integral part of the learning process.
-AMSL [38] also uses signal transformations as data augmentation.
-TimeAutoAD [18] employs three different strategies to augment the training data for generating pseudo anomalies.
-Motivated by masked language models [11], MAD [12] investigated various masking procedures for multivariateTSAD based on a predictive pretext task.
-Note that only [18, 26] parameterize the augmentation procedures, but they tune those using labeled validation data.
-Unsupervised Model Selection:
-Unsupervised hyperparameter tuning (i.e.
-model selection) is nontrivial in anomaly detection due to the absence of labeled validation data [23], and the literature is slim with quite recent efforts [39, 40, 41, 42].
-Specifically on SSL-based AD, a recent study [35] has revealed the impact of the choice of augmentation on SSL-based AD, leading to several works that aim to automatically search for the optimal choice [34, 36].
-Their main contribution is a novel unsupervised loss, but it is not differentiable.
-None of the existing works addressed the model selection problem of SSL-based AD for time series data.
+See Appx.
+D for additional results and ablation studies.
 4 Conclusion
-We introduced TSAP for self-supervised time series anomaly detection, which is the first attempt that automatically (self-)tunes the augmentation hyperparameters on time series data in an unsupervised manner.
-TSAP includes a differentiable model to augment the input time series data with various anomaly types, and an unsupervised validation loss that assists in aligning the augmented and test data.
-Experiments showed TSAP’s ability in effectively selecting the augmentation type along with its continuous hyperparameters.
-Across various real-world datasets with different types of time series anomalies, TSAP outperformed a diverse list of baselines, including modern neural and self-supervised approaches.
-While being the first self-tuning SSL solution to TSAD, our work opens new research directions, including an expanded catalog of supported anomaly types and possible extension of the ideas to multivariate time series data.
-Appendix 0.A Appendix
-0.A.1 Types of Time Series Anomalies
-We include six types of time series anomalies:
+We introduced TSAP for self-supervised time series anomaly detection, which is the first attempt that automatically (self-)tunes the augmentation hyperparameters on time series data in an unsupervised manner through a differentiable augmentation model and an unsupervised validation loss to help align augmented and test data.
+While being the first self-tuning SSL solution to TSAD, our work opens avenues for an expanded catalog of anomaly types and extensions to multivariate time series data.
+Acknowledgments
+Funding: This work was supported by the Research Foundation - Flanders (FWO) [grant number G0C6721N] and [grant number V436123N].
+Supplemental Materials
+Appendix A Types of Time Series Anomalies
+The anomaly generation scheme $g$ accommodates six types of time series anomalies as discussed in Sec.
+2.1.
+Based on $g$, the goal for $f_{\mathrm{aug}}$ is then to learn to inject anomalies into inliers, conditional on the discrete hyperparameter, anomaly type, along with their corresponding hyperparameters as described below.
+A visual overview of the anomaly types is provided in Fig.
+4.
   • Platform: Starting at timestamp location, the values of a duration length in the time series are equal to a constant value level.
@@
   • Frequency shift: Starting at phase location, the frequency of the duration with length phases is increased by a constant value level.
-Fig. 2 visualizes each type of anomaly and demonstrates how each hyperparameter controls the injection of the anomaly in the time series for platform anomaly.
-0.A.2 Dataset Details
-PhysioNet
-The 2017 PhysioNet Challenge dataset [9] comprises a diverse array of real-world 1-lead 300 Hz ECG recordings.
-We use ECG recordings, each 9 seconds in length with $K=2700$ time-steps and standardized to have a zero mean and standard deviation of one.
-Injected anomalies represent 10% of the data.
-Here, we include three additional controlled TSAD tasks based on the PhysioNet data as shown in Table 3.
-Table 4 (top) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in PhysioNet.
-The hyperparameters location and length are normalized by $K$.
-Noting that the extremum anomaly always occurs on a single timestamp in the time series, thus length is always $1$.
-Table 3: Anomaly profile of additional TSAD tasks E–G.
-Profile PhysioNet E PhysioNet F PhysioNet G
-Type Mean shift Mean shift Extremum
-Level Fixed Fixed Fixed
-Location Random Random Random
-Length Random Fixed N/A
-MoCap
-The CMU Motion Capture (MoCap) dataset [^3] http://mocap.cs.cmu.edu/ includes signal data from various sensors on subjects’ bodies as they perform different activities (walking, jumping, or running).
-As we focus on a univariate setting, only the sensor signal on the left femur is used.
-Since each data contains the time series data with a short length, we stitch them into a longer one in our experiment.
-To ensure it is done smoothly, we identified the start and end points of each gait phase for the stitching.
-We further add random noises to augment the normal samples in the dataset.
-Each signal is normalized between $-1$ and $1$ and truncated to length $K=1500$.
-Constructed anomalies represent 10% of the data.
-As opposed to PhysioNet, we consider MoCap as a dataset with natural or real anomalies.
-Therefore, we do not create additional TSAD tasks from this dataset.
-Table 4 (bottom) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in MoCap.
-The hyperparameters location and length of platform anomaly are normalized by $K$.
-The hyperparameters location and length of frequency anomaly denote the starting gait phase and the length of gait phases, respectively.
-Table 4: Hyperparameter space for each type of anomaly.
-PhysioNet
-Anomaly Types location length level
-Platform $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
-Mean Shift $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
-Amplitude $[100..2000]$ $[400..600]$ $\{0.5k+1|k\in[0..10]\}$
-Trend $[100..2000]$ $[400..600]$ $\{0.002k-0.01|k\in[0..10]\}$
-Extremum/Spike $\{100k|k\in[1..26]\}$$\{1\}$ $\{3k-15|k\in[0..10]\}$
-MoCap
-Anomaly Types location length level
-Platform $[200..800]$ $[100..200]$ $\{0.2k-1|k\in[0..10]\}$
-Frequency Shift $[1..3]$ $[1..6]$ $[1..3]$
-0.A.3 Model Configurations
-Following Sec. 4.1, we provide further details on the configurations for TSAP and all baselines.
-TSAP configuration:
-In Table 5, we provide a comprehensive overview of the configuration details for the different components of TSAP.
-Table 5 (top) shows the configuration details for the different components of $f_{\mathrm{aug}}$ following Fig.
-1 (left).
-Similarly, Table 5 (bottom) shows the configuration details for the components of $f_{\mathrm{det}}$ following Fig.
-1 (right).
+Figure 4: Examples of six different types of time series anomalies; (black) original real-world time series, (red) pseudo anomalies generated by $g$.
+The goal for $f_{\mathrm{aug}}$ is to learn to inject anomalies into inliers, conditional on the discrete (anomaly type) as well as continuous hyperparameters such as level (magnitude), location (start point) and length (see bottom right).
+Appendix B Model Configurations
+B.1 Validation Loss
+We measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space using the Wasserstein distance [10] as described in Sec.
+2.2.
+The Wasserstein distance is a distance measure between probability distributions, of order $p$ between any two marginal distributions $\mu$ and $\nu$, given by:
+$W_{p}(\mu,\nu)=\biggl{(}\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(x,y)\sim\gamma}[d(x,y)^{p}]\biggr{)}^{1/p},$ (2)
+[…]
+B.2 TSAP configuration
+The Encoderϕ in $f_{\mathrm{aug}}$ and Encoderθ in $f_{\mathrm{det}}$ are constructed using 1D CNN blocks [12] (transposed 1D CNN for Decoderϕ) for efficient temporal feature extraction.
+We carefully choose the number of epochs $T$ to allow sufficient time for the convergence of $\mathbf{a}$, with empirical evidence suggesting that $T=100$ typically suffices.
+For the number of inner-loops $L$, we opt for $L=5$, aligned with [8], such that $f_{\mathrm{det}}$ has adequate time to learn effective discriminative embeddings for $\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{trn}}$.
+Table 2 provides a comprehensive overview of the configuration details for the different components of TSAP.
 Table 2: Configuration details for $f_{\mathrm{aug}}$ and $f_{\mathrm{det}}$ of TSAP
@@
    Mixing Rate (cf. Phase ii) 0.15
-Baseline configurations:
-We used the parameter settings recommended by the original authors for all baseline models.
-This approach ensures that our comparisons are fair and align with the intended set-up of these methods originally proposed by the authors.
-  • OC-SVM: We use author-recommended hyperparameters based on [17].
-  • LOF: We use author-recommended hyperparameters based on [7].
-  • IF: We use author-recommended hyperparameters based on [21].
+B.3 Baseline configurations
+We compare TSAP with a selection of established baselines, including traditional and deep learning-based methods with demonstrated efficacy in TSAD [13].
+The traditional methods consist of different modeling approaches; namely, One-Class Support Vector Machines (OC-SVM) [14]; Local Outlier Factor (LOF) [15]; (ARIMA) [16]; Isolation Forest (IF) [17]; and the Matrix Profile (MP) [18].
+On the deep learning side, we benchmark against the Encoder-Decoder LSTM (EncDec-LSTM) [19]; the Spectral Residual Convolutional Neural Network (SR-CNN) [20]; the Unsupervised Anomaly Detection (USAD) for TSAD [21]; and a recent time series foundation model (TimeGPT) [22].
+Lastly, we include a state-of-the-art competing method which learns augmentations in the embedding space, called Neural Transformation Learning for (TS)AD (NeuTraL-AD) [7].
+This diverse set of baselines allows for a comprehensive analysis across different approaches within the TSAD domain.
+The details for the baseline configurations are provided below.
+All models were trained on a single NVIDIA Tesla P100 GPU.
+    OC-SVM: We use author-recommended hyperparameters [23].
+    LOF: We use author-recommended hyperparameters [15].
+    IF: We use author-recommended hyperparameters [17].
+    ARIMA: We use AutoARIMA to select hyperparameters [16].
+    MP: We use author-recommendations to set the window size $m$ [18]
     EncDec-LSTM: Similar to the original authors [19], we downsample our time series to obtain time series of length approx.
@@
 For the remaining hyperparameters, we use the original authors’ recommendations.
-  • SR-CNN: We use author-recommended hyperparameters based on [28].
+    SR-CNN: We use author-recommended hyperparameters [20].
     USAD: Similar to the original authors [21], we downsample our time series to obtain time series of length approx.
@@
 For the remaining hyperparameters, we use the original authors’ recommendations.
-  • NeuTraL-AD: We use author-recommended hyperparameters based on [26]. We tune augmentation type separately for each dataset using labeled validation data.
-0.A.4 Additional Results
-We present additional results for continuous and discrete augmentation hyperparameter tuning, as well as two additional ablation studies.
-0.A.4.1 A.4.1 Continuous Augmentation Hyperparameter Tuning
+    NeuTraL-AD: We use author-recommended hyperparameters [7]. We tune augmentation type separately for each dataset using labeled validation data.
+    TimeGPT: We tune the confidence interval [22] for each dataset using labeled validation data.
+Appendix C Experimental Setup
+C.1 Dataset Details
+PhysioNet
+The 2017 PhysioNet Challenge dataset [9] comprises a diverse array of real-world 1-lead 300 Hz ECG recordings.
+We use ECG recordings, each 9 seconds in length with $K=2700$ time-steps and standardized to have a zero mean and standard deviation of one.
+Injected anomalies represent 10% of the data.
+We include a total of seven controlled TSAD tasks based on the PhysioNet data as shown in Table 3, with PhysioNet A-D discussed in Sec.
+3 and the remainder in Appx.
+D.
+For example in PhysioNet A and B, given the anomaly type (Platform), the task is to infer or tune, respectively, the hyperparameter(s) level only and both level and length, while anomaly location is random (hence not tuned).
+For PhysioNet C and D, the respective tuning tasks are the same but for a different anomaly type (Trend).
+Finally, Table 4 (top) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in PhysioNet.
+The hyperparameters location and length are normalized by $K$.
+Noting that the extremum anomaly always occurs on a single timestamp in the time series, thus length is always $1$.
+Table 3: Anomaly profile of different TSAD tasks.
+Dataset Type Level Location Length
+PhysioNet PhysioNet A Platform Fixed Random Random
+PhysioNet B Platform Fixed Random Fixed
+PhysioNet C Trend Fixed Random Random
+PhysioNet D Trend Fixed Random Fixed
+PhysioNet E Mean shift Fixed Random Random
+PhysioNet F Mean shift Fixed Random Fixed
+PhysioNet G Extremum Fixed Random N/A
+MoCap MoCap A Jump Fixed Random Random
+MoCap B Run Fixed Random Random
+MoCap
+The CMU Motion Capture (MoCap) dataset [^2] http://mocap.cs.cmu.edu/ includes signal data from various sensors on subjects’ bodies as they perform different activities (walking, jumping, or running).
+As we focus on a univariate setting, only the sensor signal on the left femur is used.
+We consider the walking signal as normal data, and the signals of jumping and running as anomalies.
+To generate normal signals, we stitch the walking signals by identifying the start and end points of each gait phase and add random noise; whereas to generate anomalous ones, we stitch walking or running signals at a random location in the normal signal.
+We further add random noises to augment the normal samples in the dataset.
+Each signal is normalized between $-1$ and $1$ and truncated to length $K=1500$.
+This yields two distinct TSAD tasks as shown in Table 3.
+For each, constructed anomalies represent 10% of the data.
+Different from PhysioNet A–G where we only tune the continuous hyperparameter(s) for the given (discrete) anomaly type, for MoCap A and B, we aim to tune both the unknown anomaly type that corresponds to Jump and Run behavior, respectively, as well as the (continuous) hyperparameter level while location and length take random values.
+Finally, Table 4 (bottom) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in MoCap.
+The hyperparameters location and length of platform anomaly are normalized by $K$.
+The hyperparameters location and length of frequency anomaly denote the starting gait phase and the length of gait phases, respectively.
+Table 4: Hyperparameter space for each type of anomaly.
+PhysioNet
+Anomaly Types location length level
+Platform $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
+Mean Shift $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
+Amplitude $[100..2000]$ $[400..600]$ $\{0.5k+1|k\in[0..10]\}$
+Trend $[100..2000]$ $[400..600]$ $\{0.002k-0.01|k\in[0..10]\}$
+Extremum/Spike $\{100k|k\in[1..26]\}$$\{1\}$ $\{3k-15|k\in[0..10]\}$
+MoCap
+Anomaly Types location length level
+Platform $[200..800]$ $[100..200]$ $\{0.2k-1|k\in[0..10]\}$
+Frequency Shift $[1..3]$ $[1..6]$ $[1..3]$
+C.2 Evaluation Metrics
+Our method calculates anomaly scores on an entire sequence level $\mathbf{x}$, similar to [7].
+This is a different set-up compared to novelty detection in time series which typically operates on a point level.
+Detection on a sequence level can be especially important to spot long-term patterns (e.g.
+Trend anomalies).
+As such, we use the $F_{1}$ score and the Area Under the Receiver Operating Characteristic Curve (AUROC) as key performance metrics to quantify detection capability of anomalous sequences.
+All results are reported on the unseen $\mathcal{D}_{\mathrm{test}}$.
+We determine the optimal $F_{1}$ score, by enumerating all possible thresholds, given by the anomaly scores for a given segment $\mathbf{x}$.
+We then compute the corresponding precision and recall for each threshold and select those that yield the highest $F_{1}$.
+As such, AUROC provides a balanced view whereas $F_{1}$ shows the optimal scenario.
+Both metrics range from 0 to 1, with higher values indicating superior performance.
+Appendix D Additional Results
+In this section, we present additional results for continuous (D.1) and discrete (D.2) augmentation hyperparameter tuning, as well as several ablation studies (D.3).
+D.1 Continuous Augmentation Hyperparameter Tuning
 In addition to the results on continuous hyperparameter tuning for PhysioNet A and B (see Fig.
 3), we demonstrate TSAP’s efficacy in tuning the continuous hyperparameters on five additional TSAD tasks.
-These include PhysioNet C and D, which feature Trend anomalies (cf.
-Table 1), as well as PhysioNet E and F, showcasing Mean shift anomalies, and PhysioNet G, illustrating Extremum anomalies (cf.
+These include PhysioNet C and D, which feature Trend anomalies, as well as PhysioNet E and F, showcasing Mean shift anomalies, and PhysioNet G, containing Extremum anomalies (cf.
 Table 3).
-PhysioNet C & D:
-Figure 6: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Trend anomalies.
-Top: Given Trend anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val.
-loss (middle), and leading to high detection AUROC performance (right).
+PhysioNet C & D
+Figure 5: Tuning the continuous augmentation hyperparameter(s) with TSAP for Trend anomalies.
+Top: Given Trend anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val.
+loss (center), and leading to high detection AUROC performance (right).
 Bottom: Multiple continuous hyperparameters, here both level and length are accurately tuned to near true values (left), as guided by minimizing the val.
-loss (middle), achieving high AUROC (right).
+loss (center), achieving high AUROC (right).
 The tuning process of the continuous hyperparameters for the Trend anomalies in PhysioNet C and D is shown in Fig.
-6.
+5.
 We observe several initializations for $\mathbf{a}$ that arrive closely to the true level for PhysioNet C, as well as to the true level and length for PhysioNet D.
@@
 This shows the effectiveness of our method even in cases where anomalies are subtle.
-PhysioNet E & F:
+PhysioNet E & F
 Similarly to Trend anomalies, Mean shift anomalies are inherently subtle, especially when the level is close to zero.
-To this end, we show in Fig.
+We show in Fig.
 6 how TSAP properly tunes the continuous hyperparameters for level and length in PhysioNet E and F for several initializations.
-PhysioNet G:
+PhysioNet G
 Lastly, Fig.
@@
 Nonetheless, TSAP successfully tunes two out of four initializations and reflects – though subtly – this difference in the validation loss.
-Ultimately, this again leads to a well-tuned $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$ that performs strongly on $\mathcal{D}_{\mathrm{val}}$.
-Figure 7: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Mean shift anomalies.
-Top: Given Mean shift anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val.
-loss (middle), and leading to high detection AUROC performance (right).
+This again leads to a well-tuned $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$ that performs strongly on $\mathcal{D}_{\mathrm{val}}$.
+Figure 6: Tuning the continuous augmentation hyperparameter(s) with TSAP for Mean shift anomalies.
+Top: Given Mean shift anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val.
+loss (center), and leading to high detection AUROC performance (right).
 Bottom: Multiple continuous hyperparameters, both level and length are accurately tuned to near true values (left), as guided by minimizing the val.
-loss (middle), achieving high AUROC (right).
+loss (center), achieving high AUROC (right).
 Mean shift anomalies have low val.
 loss by default due to the subtle nature of Mean shift anomalies.
-Figure 8: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Extremum anomalies.
-Given Extremum anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val.
-loss (middle), and leading to high detection AUROC performance (right).
+Figure 7: Tuning the continuous augmentation hyperparameter(s) with TSAP for Extremum anomalies.
+Given Extremum anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val.
+loss (center), and leading to high detection AUROC performance (right).
 Note how Extremum anomalies have low val.
 loss by default due to the natural presence of spikes in ECG data.
-0.A.4.2 A.4.2   Discrete Augmentation Hyperparameter Tuning
+D.2 Discrete Augmentation Hyperparameter Tuning
 We showcased TSAP’s ability to tune the discrete hyperparameter, anomaly type, in Fig.
 3 for controlled and natural TSAD tasks.
-Given the direct applicability and significance of discrete hyperparameter tuning in real-world contexts, we present extended results for natural TSAD tasks by additionally discussing discrete hyperparameter tuning for MoCap B (cf.
-Table 1).
-MoCap B:
+Given the direct applicability and significance of discrete hyperparameter tuning in real-world contexts, we present extended results for discrete hyperparameter tuning.
+MoCap B
 Fig.
@@
 TSAP was initialized twice: first with $f_{\mathrm{aug}}$ pre-trained for injecting Frequency shift anomalies, and second with $f_{\mathrm{aug}}$ pre-trained for injecting Platform anomalies.
-The validation loss (top) indicates a strong alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and the unlabeled $\mathcal{D}_{\mathrm{val}}$ when TSAP is initialized with Frequency shift anomalies.
-This is also reflected in $f_{\mathrm{det}}$’s performance on $\mathcal{D}_{\mathrm{val}}$ (middle).
-Visually, we can indeed confirm that Frequency shift anomalies (bottom – red) appear to be more similar to the running pattern (bottom – black) as opposed to Platform anomalies (bottom – purple).
-0.A.4.3 A.4.3   Additional Ablation Studies
-Figure 9: (best in color) Tuning discrete hyperparameter (anomaly type) with TSAP.
+The validation loss (left) indicates a strong alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and the unlabeled $\mathcal{D}_{\mathrm{val}}$ when TSAP is initialized with Frequency shift anomalies.
+This is also reflected in $f_{\mathrm{det}}$’s performance on $\mathcal{D}_{\mathrm{val}}$ (center).
+Visually, we can indeed confirm that Frequency shift anomalies (right – red) appear to be more similar to the running pattern (right – black) as opposed to Platform anomalies (right – purple).
+Figure 8: Tuning discrete hyperparameter (anomaly type) with TSAP.
 For Run anomalies in MoCap B with unknown type (black), val.
-loss favors Frequency shift (red) that leads to high AUROC (middle) and mimics well true anomaly (bottom), and effectively rejects type Platform (purple) with poor performance.
-Following Sec.
-4.4, we present additional ablation studies for the second-order optimization within TSAP and the normalization of the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ obtained through $f_{\mathrm{det}}^{\mathrm{enc}}$.
-Second-order Opt. Ablation:
+loss favors Frequency shift (red) that leads to high AUROC (center) and mimics well true anomaly (right), and effectively rejects inferior type Platform (purple).
+D.3 Ablation Studies
+We present four ablation studies on the controlled PhysioNet data to support various design strategies of TSAP: validation loss, self-tuning, second-order optimization, and embedding normalization.
+Validation Loss
+In Fig.
+9 (top), we illustrate the level estimation for PhysioNet C under the condition where our Wasserstein-based $\mathcal{L}_{\mathrm{val}}$ is substituted with a point-wise metric, as used in [8].
+This comparison shows that a point-wise validation loss tends to favor solutions where the level of the Trend anomaly approximates zero, essentially neutralizing the anomaly.
+Although this might produce high alignment, it leads to poor $f_{\mathrm{det}}$ performance in $\mathcal{D}_{\mathrm{val}}$ (right).
+This shows that the distributional characteristics captured by our $\mathcal{L}_{\mathrm{val}}$ are a key contributing factor to the success of TSAP.
+Figure 9: Overview of ablation studies.
+Top: TSAP’s $\mathcal{L}_{\mathrm{val}}$ is replaced by a point-wise val.
+loss leading to an erroneous estimation of $a$ (left) and poor performance (right).
+Bottom: TSAP’s self-tuning module is disabled, $a$ is now randomized.
+Val.
+loss indicates poor alignment, reflected in poor performance.
+Randomization vs. Tuning
+In Fig.
+9 (bottom), the self-tuning module is disabled for PhysioNet C, where level is instead randomized (along with location, and length).
+We observe substantially higher $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment.
+In turn, $f_{\mathrm{det}}$ struggles to detect the unlabeled anomalies in $\mathcal{D}_{\mathrm{val}}$, showing the utility of TSAP’s systematic hyperparameter (self-)tuning over random choice.
+Second-order Opt. Ablation
 Fig.
@@
 In turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ suffers severely.
-Embedding Normalization Ablation:
-We show the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when embedding normalization is disabled in Fig.
+Embedding Normalization Ablation
+We show the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when the normalization of the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ obtained through $f_{\mathrm{det}}^{\mathrm{enc}}$ is disabled in Fig.
 10 (bottom).
-While $a$ initialized at $0.01$ eventually leads to the correct level, the estimation process is much more volatile compared to when normalization is enabled as shown in Fig.
+While $a$ initialized at $0.01$ eventually leads to the correct level, the estimation process is highly volatile compared to when normalization is enabled as shown in Fig.
 5 (top).
-Figure 10: (best in color) Overview of additional ablation studies.
+Figure 10: Overview of additional ablation studies.
 Top: We disable second-order optimization in TSAP, leading to a highly unstable estimation process of $a$ (left) and poor performance of $f_{\mathrm{det}}$ (right).